### v1

# Blockchain Address Poisoning

Abstract
In many blockchains, e.g., Ethereum, Binance Smart Chain (BSC), the primary representation used for wallet addresses is a hardly memorable 40-digit hexadecimal string. As a result, users often select addresses from their recent transaction history, which enables blockchain address poisoning. The adversary first generates lookalike addresses similar to one with which the victim has previously interacted, and then engages with the victim to “poison” their transaction history. The goal is to have the victim mistakenly send tokens to the lookalike address, as opposed to the intended recipient. Compared to contemporary studies, this paper provides four notable contributions. First, we develop a detection system and perform measurements over two years on Ethereum and BSC. We identify 13 times the number of attack attempts reported previously—totaling 270M on-chain attacks targeting 17M victims. 6,633 incidents have caused at least 83.8M USD in losses, which makes blockchain address poisoning one of the largest cryptocurrency phishing schemes observed in the wild. Second, we analyze a few large attack entities using improved clustering techniques, and model attacker profitability and competition. Third, we reveal attack strategies—targeted populations, success conditions (address similarity, timing), and cross-chain attacks. Fourth, we mathematically define and simulate the lookalike address-generation process across various software- and hardware-based implementations, and identify a large-scale attacker group that appears to use GPUs. We also discuss defensive countermeasures.
1 Introduction
(a) Tiny transfer
(b) Zero-value transfer
(c) Counterfeit token transfer
Figure 1: Poisoning transfers
Most modern blockchains rely on “wallets” for monetary transfers. Wallet addresses are derived from cryptographic public keys and are often represented by long, hard to memorize, strings. However, most currency or token transfers require users to manually input or select the recipient’s wallet address. A common practice is to copy and paste addresses or select an address with which one has previously interacted. This introduces a new attack, blockchain address poisoning.
An adversary generates one or more “lookalike” address(es) whose first and last characters match those of an address the victim often interacts with. The attacker then floods (“poisons”) the target’s transaction history with the lookalike address(es). As a result, the victim may erroneously send digital assets to a lookalike address instead of the intended address. In contrast to traditional bank transfers, blockchain transactions are irreversible, making fund recovery challenging, and thereby exacerbating the severity of any mistakes. In this paper, we attempt to characterize the economics of address poisoning and to uncover attackers’ strategies and capabilities through large-scale measurements and simulations.
We design a detection system and run it between July 1, 2022, and June 30, 2024, on Ethereum and Binance Smart Chain (BSC). Compared to contemporary studies [46, 10], our detection 1) covers all types of poisoning attacks, 2) demonstrates greater robustness against attacker manipulation, and 3) generalizes better across time intervals and chains.
Combining the results from both chains, we identify over 270M attack attempts (i.e., 13 times the previous efforts: 21M [46] and 14M [10]) that target over 17M victims. Out of those, attackers successfully received 6,633 transfers, adding up to over 83.8M USD in ill-gotten gains. This attack appears to be one of the largest cryptocurrency phishing schemes observed in the wild. We validate our results by 1) manually verifying the phished cases, 2) directly comparing the results with the related works, 3) checking the smart contracts used for those attacks, and 4) comparing the data collection results with different parameters (e.g., a window size).
Of independent interest, we leverage our detection system to capture accidental transfers (e.g., typing mistakes in the recipient’s address). We confirm 681 accidental transfers with 5.5M USD losses, based solely on human mistakes which underscores the need for more forgiving user interfaces.
Attackers typically optimize their operations by bundling several transfers into a single transaction and reusing both addresses and contracts. This behavior allows us to link attack addresses through the “guilt-by-association” heuristic [43]. After removing transactions from unrelated bots that occasionally copy poisoning transactions from attackers, we can establish the existence of a few large distinct attack groups which we use to model the economics of address poisoning. For each group, we calculate revenues (i.e., a sum of phished amounts) and costs (i.e., total attack transaction costs). Despite high variability, most large groups manage to stay profitable by succeeding often, while smaller groups sometimes suffer from loss (by losing to the competition). There is no apparent early mover advantage.
We also analyze attack strategies. First, we evaluate the characteristics of the targeted victims to learn what attackers are looking for. We find that targets are more likely to have had higher balances at the time of the attack, conducted more transactions, and transferred a larger amount compared to random stablecoin users. Second, we investigate what influences attack success. We find that victims are more likely to choose a lookalike address that is more similar to or appears earlier than others. Third, given that addresses are compatible across Ethereum and BSC, some attack groups re-use lookalike addresses and target victims across chains.
To infer attackers’ (hardware) lookalike address generation capabilities, we simulate address generation rates. We implement our scripts both naively and with optimizations and compare the performance over different CPUs and GPUs. Based on our benchmarks, one sophisticated attack group appears to be using GPUs while others seem to rely on CPUs.
Finally, we show that attackers are not as efficient as they could be, suggesting that the attack would be even more powerful if they optimized their strategies. This motivates us to propose a few mitigations for address poisoning attacks at the protocol, contract, wallet, and user levels.
2 Background
We next discuss 1) basic blockchain primitives (smart contracts, ERC-20 tokens), 2) wallet types and addresses, and 3) differences between address poisoning and phishing attacks.
2.1 Blockchain primitives
Most popular modern public blockchains are either UTXO (Unspent Transaction Output)-based or account-based. On UTXO-based chains such as Bitcoin, address reuse is minimal: most addresses are used only for two transactions (one to receive some money, one to spend it). On the other hand, account-based chains such as Ethereum see significant address reuse. Account-based chains are therefore acutely susceptible to address poisoning as users frequently reference recipient addresses from past transactions. In this paper, we only focus on account-based chains, so we use “account” and “wallet addresses” interchangeably.
Ethereum (and derived blockchains) features two types of addresses: wallet addresses (also called External Owned Accounts—EOA) and contract addresses. Wallet addresses to send Ether (ETH, the native currency) or interact with smart contracts (whose addresses are contract addresses). Smart contracts are executed by all nodes using the Ethereum Virtual Machine (EVM). Smart contracts are developed in a higher level programming language (e.g., Solidity), and then compiled to “bytecode” intelligible by the EVM before being deployed. Smart contracts allow users to perform actions, such as token implementation, token transfer. Of particular interest, ERC-20 tokens (e.g., USDT, USDC) are fungible assets, defined by the ERC-20 framework [40], for which creators specify a token name, symbol, and value unit. Essentially ERC-20 tokens allow developers to implement other cryptocurrencies or payment systems on top of Ethereum.
For every action (e.g., sending ETH or ERC-20 tokens, or more generally, interacting with any contract), users pay “gas,” that is, transaction fees dependent on the computational complexity of the action requested. Other chains, such as Binance Smart Chain (BSC), that inherit Ethereum’s structure are called EVM-compatible chains. Notably, wallet addresses are compatible across EVM chains, allowing users to hold tokens across chains in one wallet.
2.2 Wallet and blockchain address
Most users will interface with the blockchain via wallet software such as MetaMask or Phantom. These powerful tools enable users to have full control over their funds and perform complex tasks such as engaging in DeFi (decentralized finance) activities. When sending direct payments, users must often manually specify the recipient address, exacerbating usability challenges—since addresses are represented by long (40-character), hardly memorable, hexadecimal strings. This UX challenge facilitates a significant threat from phishing or scams and motivates several concerns about wallet interfaces (e.g., the absence of strong enough warnings) [47].
Generating a wallet address in an EVM-compatible chain is a three-step process:
(1) Choose a 32-byte private key ($k$), which must be truly random to avoid a wallet breach [1].
(2) Apply elliptic curve (ECDSA) multiplication to produce a public key $K=k\times G$ where $G$ is the fixed generator point on the elliptic curve.
(3) Hash (Keccak 256) the public key $K$, take the last 20 bytes, and add prefix $0x$ to obtain the address.
Elliptic curve (ECDSA) multiplication is computationally hard to invert, meaning that it is impractical to derive a private key from a public key. So, to generate a public key with a certain format (e.g., a vanity address), the only viable strategy is to perform a brute-force computation over potentially many private keys. The complexity of the brute force computation depends on the number of characters that have to be fixed in the desired public key address. For instance, it is roughly $16^{3}$ times harder to generate an (hexadecimal) address with a 3-character desired prefix compared to a random public key address. Attackers adopt this brute-force technique to generate lookalike addresses.
2.3 Attack characteristics
Phishing attacks on domain names have several variants. Examples include typosquatting [21], which involves users mistyping domain names; combosquatting [16], wherein attackers append related keywords to domain names; and homograph attacks [7], where attackers substitute characters in a domain name with visually similar ones (e.g., Cyrillic characters instead of Latin [13]). All these attacks exploit human features—our inability to distinguish between graphically near-identical representations, as well as our ability to quickly make inferences from incorrect text [28]. Domain-name-based phishing attacks may resort to indirect monetization, e.g., ad impressions, or affiliate program enrollment [13, 21, 2].
Blockchain address poisoning is markedly different from domain-name based phishing attacks in several respects. First, it focuses on the direct extraction of funds, and it may be more challenging for a victim to recover funds due to the irreversibility of cryptocurrency payments.
Second, blockchain address poisoning attempts are permanently recorded in a public ledger, allowing for more accurate estimates of the losses [19]. Estimating the impact of traditional phishing attacks is challenging due to data confidentiality [24] and phishing website ephemerality [39, 37, 3, 19].
As a third difference, address poisoning targets hexadecimal strings, while domain-name based attacks typically focus on character strings. Previous typosquatting studies introduced various metrics for quantifying domain name similarity, such as Damerau-Levenshtein (DL) distance [6], fat finger (FF) distance [21], and visual distance [32]. However, blockchain address poisoning involves copy-and-paste or multiple-choice selection rather than manual typing of characters. Additionally, long hexadecimal addresses are often abbreviated to their prefix and suffix, especially on constrained interfaces such as mobile. Thus, DL and FF distances are less relevant here. We instead use the number of matching hexadecimal digits at the beginning and at the end.
Fourth, domain-name based attacks often proceed in multiple stages to steal funds: the victim first clicks on a malicious URL, is directed to a phishing website and is prompted to input sensitive information that can eventually be monetized. On the other hand, blockchain address poisoning is far more direct. Victims only need to make the single mistake of submitting transactions by selecting the attacker’s address.
Finally, blockchain address poisoning has differing costs to the attacker. They only need to deploy attack smart contracts (which can be even recycled from others for free), and one poisoning transfer costs only about a dollar in Ethereum and a cent in BSC. Generating lookalike addresses however has a variable cost that can be adjusted by the attacker, ranging anywhere from a laptop CPU to data centers of GPUs. Domain-based phishing attacks, on the other hand, require registering/hijacking domains, crafting phishing websites, and/or buying tool kits [3], all of which is expensive and time-intensive.
3 The attack
This section describes address poisoning in detail. We first introduce our threat model and attacker capabilities, discuss various ways to carry out address poisoning, and examine how attackers execute them in practice.
3.1 Threat model
We assume the existence of an attacker that passively monitors all public blockchains simultaneously in real-time. When available the attacker may also observe a state that has yet to be finalized in a chain such as transactions in the mempool. The attacker can also actively engage with any chain by launching contracts and sending transactions using many accounts, bounded only by the chain and economic constraints. Finally, the attacker can generate lookalike addresses at a cost consistent with (publicly known) state-of-the-art hardware.
We do not assume that the attacker has any out-of-band information about or agency towards a victim other than what can be obtained through public blockchains. Therefore attackers cannot influence the wallet software a victim is using, or make predictions about a victim’s transfers beyond what can be inferred through public signals. We also assume that the security models for the individual chains hold, so attackers cannot, for example, perform denial of service on a chain or drop/modify a victim’s transactions.
The attacker’s ultimate goal is to have the victim send assets to (one of) the attacker’s lookalike address(es) instead of the victim’s intended address. To do so, the attacker first monitors blockchain transactions and identifies asset (ETH or ERC-20 tokens) transfer events. When such transfers are observed, the attacker identifies an intended recipient wallet, $R$, and generates a lookalike address $L$. We will later numerically evaluate the computational requirements to get lookalike addresses convincing enough for the attack to be viable.
The attacker then submits to the chain phishing transactions that interact with the victim $V$ and contain the lookalike address $L$, typically within a short period of time. When the victim later attempts to send assets to $R$, they could get confused and mistakenly send funds to $L$ instead. This confusion typically arises through some combination of UI/UX limitations, and the victim failing to carefully check the recipient. A common vector is for a victim to transfer funds to an address that is pre-populated by their wallet software in a list of addresses with which they had recent interactions. We focus on ERC-20 (or BEP-2 in BSC) tokens for the underlying asset as attacks with these tokens are most prevalent.
3.2 Transfer types
The attacker poisons the victim through three types of token transfers: tiny transfers, zero-value transfers, and counterfeit token transfers. Figure 1 depicts these variations.
In a tiny transfer (Fig. 1(a)), the attacker always sends from $L$ to $V$ a small amount of the same token that $V$ had previously sent to $R$. The idea is that $L$ now shows up in $V$’s recent transaction history. Later, $V$ could mistakenly assign $L$ as the recipient due to its similarity to $R$. This attack is similar to blockchain dusting [42], in which the attacker sends many small transactions to a victim for deanonymization.
In a zero-value transfer (Fig. 1(b)), the attacker creates a transfer from $V$ to $L$ for the same token that $V$ sent to $R$, but with the value of zero. Typically, token transfers can only be sender-initiated. However, the current token standard features a public function called transferFrom, and popular token implementations allow the function callers to specify an arbitrary sender and an arbitrary receiver if the transferred value is zero. Therefore, the attacker can record a transfer event from $V$ to $L$ of the same ERC-20 token with zero value. Zero-value transfers are potentially even more dangerous than tiny transfers, as the victim appears to have previously sent money to the lookalike address.
In a counterfeit token transfer (Fig. 1(c)), the attacker sends from $V$ to $L$ the same amount $V$ sent to $R$, but with a counterfeit token created by the attacker (“USDTT” here). Since the attacker controls the implementation of the counterfeit token, they can have the ability to transfer it on behalf of anyone. There is no restriction on names, symbols, or functionality for ERC-20 tokens [8, 43], which makes it easy to create similar-looking tokens for malicious purposes. Because most popular wallet software maintains a curated list of popular (chain, contract address, symbol) tuples, attackers often avoid names (or symbols) identical to mainstream tokens to avoid triggering a warning. In the case of new or recently launched tokens, they could even proactively register the token name first [46].
Finally, these types of transactions are not mutually exclusive. Indeed, the attacker can combine these strategies, flood the victim with transactions, and wait for a mistake.
3.3 Attack implementation
Figure 2: Attack mechanism
Figure 2 shows one method of implementing blockchain address poisoning in practice. Wallets are represented by rectangles and contracts (or ERC-20 tokens created by contracts) are represented by ovals. We color attackers’s instances in pink. First, an attack account ($A$) deploys a tailor-made attack contract ($AC$) and calls a function from it to execute multiple poisoning transfers at once. The attacker can also directly interact with the token contract to initiate a single poisoning transfer (without $AC$), which we observe in some cases despite its higher gas cost. The transaction $TX$ bundles together several potentially unrelated phishing transfers, in this case, the transfer of a lookalike token from $V1$ to $L1$ and a zero-value transfer from $V2$ to $L2$. Notice also that attacker $A1$ recycles their contract $AC1$ and $CT1$ to perform additional phishing transfers in $TX3$, and the infrastructure that $A1$ has embedded on a chain is typically distinct from that of other attackers such as $A2$’s use of $AC2$. It is also common for multiple attackers to target the same victim (e.g., $V1=V4$ ). As we will later describe, we have observed cases where attackers launch more than 100 poisoning transfers in a single transaction.
4 Detecting attacks
We next explain how we implement our detection system and how we cluster individual attack instances into attack groups.
4.1 Detection algorithm
Figure 3: Detection algorithm sketch
Ye et al. [46] identify zero-value transfers by enumerating all token transfers with zero value. While computationally efficient, this method does not consider tiny and counterfeit token transfers and possibly mislabel benign zero-value transfer attacks, e.g., those used for testing a destination address. Guan and Li [10] capture all three poisoning transfers but follow an approach slightly different from ours. They flag suspicious transfers based on attack characteristics first, whereas we filter transfers when a victim interacts with two improbably similar, but different addresses. Our approach appears more robust to attacker manipulation (refer to Appendix A). We deploy our own Ethereum and BSC full nodes to scan the data efficiently instead of relying on third-party RPC endpoints. Figure 3 sketches our detection system, which consists of five steps.
Step 1: Identifying potential victims. For block $n$, we collect the list of addresses (senders) $\mathcal{V}=\{V_{1},V_{2},\ldots\}$ who have sent any amount of one (or more) of the major stablecoins [^1] Tether (USDT), USD Coin (USDC), and DAI for Ethereum and Binance-Peg USD Coin (USDC), Binance-Peg BSC-USD (BSC-USD), Binance-Peg BUSD (BUSD), and any USDC for BSC. and the addresses =$\mathcal{R}=\{R_{1},R_{2},\ldots\}$ that have received these tokens. $\mathcal{V}$ is the set of (potential) victim addresses, and $\mathcal{R}$ is the set of intended addresses.
Step 2: Identifying possible poisoning transfers. We next collect potential poisoning transfers, i.e., the attempt(s) to tamper with the victim’s address lists. These poisoning transfers often happen soon after the attacker has identified a potential victim through the original token transfer (which we collected in Step 1). So, we focus on the 20 minutes following the original transfer, which corresponds to blocks $n+1$ to $n+m+1$ where $m=100$ and $400$ for Ethereum and BSC, respectively. We validate this choice of window size in §5.2.
We look at all transfers in stablecoins – i.e., we search all transfers sent to (or from) addresses in $\mathcal{V}$ for potential tiny (or zero value) transfers. We also retrieve all ERC-20 token transfers for which an address in $\mathcal{V}$ is a source for potential counterfeit token transfers. We denote the set of recently used addresses that interact with $\mathcal{V}$ by $\mathcal{C}=\{C_{1},C_{2},\ldots\}$.
We iterate this two-step process for every block $n$ over a two-year interval. [^2] We collect transfer events through the eth_getLogs JSON-RPC API with signature 0xddf252ad1be2c89b69c2b068fc378daa952ba7f163c4a11628f55a4df523b3ef. Given that attackers often launch multiple poisoning transfers within a single transaction, we further retrieve other transfer events (if any) in the same poisoning transactions. For instance, in Figure 2, if we detect $TR_{1}$, we can also capture $TR_{2}$.
[…]
Step 4: Pruning possible attacks. For each victim $V$, the corresponding intended address $R$, and the associated lookalike address $L$, we extract all related token transfer events collected in step 2. In other words, we only look at the cases where the victim $V$ interacts with two very similar addresses $R$ and $L$, making false positives statistically unlikely (except for those caused by typing mistakes, which we separately discuss in Appendix C).
To identify counterfeit tokens, we acquire a list of ERC-20 and BEP-20 tokens from two sources: 1) Etherscan and BSCscan [^4] Tokens with at least a “Neutral” or “OK” reputation from Etherscan: https://etherscan.io/tokens and BSCscan: https://bscscan.com/tokens as of August 13, 2024. and 2) Coinranking. [^5] https://coinranking.com/coins/erc-20 as of August 31, 2024. The union of these lists forms the set of authentic tokens $\mathcal{AT}$; all others form the set of counterfeit tokens $\mathcal{CT}$.
We then categorize the transfers we have collected into three major categories (intended, poisoning, payoff), distinguishing poisoning transfers between tiny, zero-value, and counterfeit token transfers per their description in Figure 1. Let $v$ be the monetary value of the transfer. Formally:
  • A transfer is an intended transfer if a potential victim $V$ sends an authentic token $AT\in\mathcal{AT}$ to an intended recipient $R$, and $v>0$.
  • A transfer is a poisoning transfer:
    1. 1.
      If a lookalike address $L$ sends an authentic token $AT$ to a victim $V$ with $0<v<v_{0}$ (tiny transfer attack). We set the threshold $v_{0}$ to 10 USD to exclude unreasonably high transfers from the attacker. [^6] More than 98.79% of tiny transfers below 10 USD are less than 3 USD. We manually found some rare cases where the attackers sent nearly 10 USD as a tiny transfer.
    2. 2.
      Or, if a victim $V$ sends an authentic token $AT$ to a lookalike address $L$ with $v=0$ (zero-value transfer attack).
    3. 3.
      Or, if a victim $V$ sends a counterfeit token $CT\in\mathcal{CT}$ to a receiver $L$ (counterfeit token transfer attack). Recall that because the token was created by the attacker, they could implement any semantics they desire, notably having anybody send tokens to anybody else, without any authentication.
  • A transfer is a payoff transfer if the victim $V$ sends an authentic token $AT$ ($v>0$) transfer to a lookalike address $L$. This is the transfer that ultimately monetizes the attack for the adversary.
The top-right table in Figure 3 summarizes our labeling scheme. Red cells are attackers. Note that we convert the transferred value into USD on the day of the transfer. The price data is from CoinGecko API based on a contract address.
Step 5: Verifying payoffs. Among all possible payoff transfers found in Step 4, we consider a transfer to be confirmed if $V$ has received at least one poisoning transfer from $L$ between the intended transfer and the (candidate) payoff transfer.
Others are unconfirmed payoff transfers—the victim sends $AT$ to $L$ without having received a poisoning transfer from $L$. This can happen for three reasons.
First, the victim could have legitimately interacted with two similar addresses by sheer coincidence. This is statistically unlikely, specifically because we already filter out addresses particularly vulnerable to the “birthday paradox” (refer to Appendix B for our detailed analysis).
Second, the attack may have taken place over a time interval exceeding our 20-minute monitoring window. To address this, for each unconfirmed payoff transfer, we fetch the victim’s entire token transfer history (from Etherscan and BSCscan) between the intended transfer and the unconfirmed payoff transfer. If we find any poisoning transfers we had failed to capture, we confirm the payoff transfer.
Third, the victim could have accidentally sent tokens to an incorrect destination address, e.g., due to a typing mistake. If 1) the destination address never sent any tokens, i.e., the attacker does not redeem the payoff transfer, and 2) $d$ is more than 20, i.e., computationally infeasible to generate such a $L$, we assume that the destination address was actually a typo. We indeed find 681 accidental transfers from 501 users, totaling 5.5 million USD in losses (almost all in Ethereum, with only 57,000 USD in BSC): Appendix C for more detailed method, results, and analysis.
4.2 Clustering attackers
Following traditional cybercrime literature [33], we want to link attack instances to extrapolate attack groups and analyze strategies, profitability, and infrastructure. Each address poisoning attack consists of 1) a poisoning transfer ($TR$), 2) a poisoning transaction ($TX$) that initiates $TR$, 3) a lookalike wallet address ($L$), 4) a counterfeit token ($CT$), 5) an attack contract ($AC$) that initiates $TX$, and 6) an attack wallet address ($A$) that interacts with $AC$ (see Figure 2).
We associate each poisoning transfer $TR$ with the corresponding attack instance $\{TX,L,CT,AC,A\}$. Similar to the “guilt-by-association” technique [43], we merge two attack sets if they share at least one of $TX$, $L$, or $A$. In other words, we assume that 1) two poisoning transfers in the same transaction belong to the same attacker, 2) an attacker is economically rational and specifies addresses $L$ they control (to successfully receive payoffs), and 3) an attacker does not share the private key of the attacker’s wallet $A$ with other attackers. We do not use information on $AC$ or $CT$ to avoid spurious cluster merges because anyone can call an attack contract or use a counterfeit token, regardless of source code availability. For instance, in Figure 2, we 1) merge ($TR_{1}$, $TR_{2}$) because both transfers are in the same transaction, 2) merge ($TR_{1}$, $TR_{3}$) if $L_{1}=L_{3}$, and 3) merge ($TR_{3}$, $TR_{4}$, $TR_{5}$, $TR_{6}$) if $A_{1}=A_{2}$.
Having applied this process to our entire dataset, we notice that two seemingly distinct large attack groups—active at different times or using different attack strategies—can be merged due to a few overlapping attack addresses $A$. Manual investigation revealed that those accounts appear to be bots copying poisoning transactions from two different attack groups with little or no modification, leading to a seemingly erroneous merge. Indeed, Guan and Li [10] performed similar clustering techniques, but did not consider the existence of bots and appeared to form one large cluster.
To prevent such issues, we explored several methods to exclude transactions from bots. While most of the approaches either still produce erroneous merges or remove too many transactions and underestimate the group size, analyzing account history effectively addresses both issues. For each attack address $A$, we additionally fetch their account history from Etherscan and calculate an attack ratio, defined as the proportion of transactions used for poisoning transfers. We then set a threshold to exclude addresses that have a low attack ratio. We document the details of those bots’ behavior and the bot exclusion process in Appendix D.
5 Results and evaluation
We apply our detection algorithm over the period Jul. 1st, 2022 to Jun. 30th, 2024 for both Ethereum (blocks 15,053,226 to 20,207,948) and BSC (blocks 19,170,674 to 40,077,592). The chain scanning process (written in Go) takes over two weeks on Ethereum and over a month on BSC (in Ubuntu 20.04.5 LTS server; AMD EPYC 9124 16-Core Processor; 128GB RAM), highlighting the need for our optimizations.
5.1 Detection results
We present the summary statistics for poisoning transfers in Table 1 and confirmed payoff transfers in Table 2
Table 1: Attack summary statistics
Blocks Transactions All poisoning Tiny Zero-value Counterfeit-token Victim Lookalike Attack Counterfeit transfers transfers transfers transfers addresses addresses contracts tokens Ethereum 5,154,722 1,691,529 17,365,954 308,881 7,185,298 9,871,775 1,330,948 6,492,215 3,480 6,280 BSC 20,906,918 16,505,215 252,703,515 3,651,015 140,556,905 108,495,595 16,107,774 43,644,433 406 710 Total 26,061,640 18,196,744 270,069,469 3,959,896 147,742,203 118,367,370 17,438,722 50,136,648 3,892 6,990
Table 2: Successful phishing attacks
Total Loss Min Median Avg. Max Std. dev Count Victims (USD) (USD) (USD) (USD) (USD) (USD) Ethereum 79,344,412 0.001 2,160 45,653 20,000,000 521,663 1,738 1,502 BSC 4,490,804 0.000 85 1,164 279,489 8,366 4,895 4,004
On Ethereum, we find 17.3M poisoning transfers: 300,000 tiny transfers, 7.2M zero-value transfers, and 9.9M counterfeit-token transfers over 1.7M transactions. Attackers target 1.3M victim addresses from 6.5M lookalike addresses. On Feb. 18th, 2023, we capture 362,934 poisoning transfers ($\approx$50 transfers per block). Those attack transactions consume 6.6% of total gas used that day, producing wasteful state on the chain and increasing network operating cost [^7] We assume the average block gas limit of 15 million, and aggregate gas usage of all poisoning transactions to calculate the percentage..
Attackers on Ethereum receive 1,738 transfers from 1,502 victims, totaling nearly 80M USD (avg. 45,853 USD, median 2,169 USD, with high variance), with a success rate is 0.01%8880.13% spam click-through rate in Twitter [9]. Some victims sent assets to the attacker’s address(es) multiple times. For instance, one victim sent 1.999M and 2M USDC within 10 blocks without noticing they were being phished.
On BSC, the number of attacks is significantly larger. We identify over 252M poisoning transfers: 3.6M tiny transfers, 141M zero-value transfers, and 108M counterfeit-token transfers in 17M transactions. Attackers target 16M victim addresses from 44M lookalike addresses. On Jun. 5th, 2024, we find over 3M poisoning transfers (i.e., $\approx$105 transfers per block). The result suggests a higher attack prevalence in chains with lower transaction fees, leading to a significant clutter in UIs and degradation in user experience. Attackers on BSC receive 4,895 transfers from 4,004 victims, totaling 4.5M USD (avg. 1,164 USD, median 85 USD).
As discussed in §4, we only focus on a few stablecoins for data collection efficiency, so our estimate is a conservative lower bound. For instance, in a well-known address poisoning attack, a victim mistakenly sent 1,155 WBTC ($\approx$68M USD) to a lookalike address in Ethereum [34, 38]. We capture the attack attempt (which belongs to the 13th largest group in our clustering results in Section 5.3) but we do not capture the final transfer (i.e., WBTC which is not a stablecoin), so our loss estimates do not include this attack.
5.2 Detection evaluation
Manual evaluation on successful cases. Statistically, the probability of getting a false positive is exceedingly low because 1) we only look at transfers where the victim has interacted with two similar addresses, and 2) we observe poisoning transfers between an intended and a payoff transfer take place. Nevertheless, two of the authors also manually verified transfers labeled as successful payoff transfers. We sorted payoff transfers by decreasing the amount and checked the 30 largest transfers for both chains. We visit each victim’s page on external blockchain scanning websites (e.g., Etherscan and BSCscan) and manually confirmed 1) the intended (original) transfer, 2) the poisoning transfer, and 3) the payoff transfer. We verified that all 30 cases are indeed payoff transfers for both Ethereum and BSC (i.e., no false positives). Those 30 cases account for 53.5M (67%) of Ethereum losses and 2M (45%) of BSC losses. For accountability purposes, we list the top 30 successful cases for Ethereum and BSC in Appendix G.
Comparison with previous work. We directly compare our detection results by focusing on 12 and 16 months of data to align with Ye et al.’s work [46] and Guan and Li’s [10] work, respectively. We detect more payoff transfers than Ye et al. We find 1,059 successful cases with 36.6M USD in losses while Ye et al. found 560 successful cases with at most 28M USD in losses, including two separate (non-address poisoning) attacks they investigate. In terms of poisoning transfers, we identify 4.9M zero-value token transfers, compared to 21M for Ye et al. It is unclear whether the difference comes from 1) false positives on Ye et al.’s side, (i.e., benign zero-value token transfers), or 2) false negatives on our side.
We find a similar total loss on Ethereum (during the same period) as Guan and Li [10], but we detect more than 2.2M counterfeit token transfers, mostly from tokens with modified symbols (refer to a more detailed comparison in Appendix A).
Checking the contract verification. We may mislabel counterfeit token transfer attacks if a transferred token is authentic but not in our ERC-20 token list ($\mathcal{AT}$). In general, public ERC-20 token creators upload the smart contract source code to Etherscan, get verified by Etherscan, and allow users to interact with (e.g., call a function of) that contract. Attackers, on the other hand, have few incentives to upload their source code. By leveraging this fact, we check the existence of the smart contract source code (i.e., verification) on Etherscan for each counterfeit token $CT$. We also examine each attack contract $AC$ for the same reason. 7 (out of 6,280) counterfeit token contracts and 8 (out of 3,480) attack contracts are verified on Etherscan. These involve only 14 out of 17.3M poisoning transfers. Such potential false positives could happen when the victim specifies a similar address (by a typing mistake or by coincidence) and uses a very minor (but not counterfeit) token (i.e., $\notin\mathcal{AT}$) at the same time. Some of those verified contracts are false positives (e.g., authentic tokens that are not in our dataset), but others are indeed counterfeit contracts that attackers uploaded to Etherscan. In other words, some attackers had uploaded the source code of the attack contracts—incidentally, this helped us illuminate how the attack is conducted (which we described in §3). We exclude verified attack contracts and counterfeit token contracts from our clustering to avoid potential erroneous merge.
Window range. To make the chain scanning process efficient, we assume that the attack happens in the next 20 minutes (100 blocks in Ethereum) after the original transfers. We evaluate this assumption by collecting data over 200,000 blocks starting from 16,950,603 (Apr. 1st, 2023), using both a 100-block range and a 200-block range, and comparing results. Both ranges identify the same number of successful cases (90). The 200-block range only identifies two additional tiny transfers besides the 250 tiny transfers identified with a 100-block range. With a 100-block range, we find 102,154 zero-value transfers; extending to 200 blocks only identifies an additional 27 zero transfers. Finally, a 100-block window captures 704,998 counterfeit token transfers, whereas a 200-block window brings this up to 706,332. In short, doubling the window size (20 minutes to 40 minutes) does not dramatically increase the number of attacks we discover. We believe that collecting other transfers within the same poisoning transaction (§4.1: Step 2) has significantly helped the coverage.
5.3 Clustering results
Before clustering, we attempt to remove addresses (bots) that copy transactions by setting an attack ratio threshold, as discussed in §4.2. We choose a threshold of 0.5, which 1) avoids erroneous merge of large clusters and 2) only removes 1.49% of all transfers. We provide a detailed explanation of this threshold selection and demonstrate the robustness of our clustering results (through temporal clustering) in Appendix D.
Table 3 describes the number of attack instances and assigns Group IDs sorted by the number of lookalike addresses in each group on Ethereum. The notation follows §3. We only list the top eight biggest groups because the size of the eighth group is significantly smaller than the top seven (i.e., less than 100,000 lookalike addresses with a few successful cases). Different groups exhibit different attack strategies; not all groups use three poisoning transfers. Only Group 1 uses tiny transfers. Groups 3, 4, and 7 do not have any counterfeit token contracts, $CT$, and thus do not use any counterfeit-token transfers, and instead focus on zero-value transfers.
We additionally collect contract bytecode from our Ethereum node and list the number of distinct contracts in a parenthesis. In Group 2, for example, all $CT$ and $AC$ are identical, indicating the reuse of the same source code (but with different addresses). Given that the contract bytecode changes even with a slight alteration in the source code, the high number of identical contracts corroborates the validity of our clustering techniques (especially since our clustering does not rely on those contracts).
Table 3: Attack group general statistics. Notations follow Fig. 2). The distinct number of contracts is in parentheses.
Group $\mathbf{L}$ $\mathbf{CT}$ $\mathbf{A}$ $\mathbf{AC}$ $\mathbf{R}$ $\mathbf{TR}$ $\mathbf{TX}$ 1 1,390,902 3,309 (517) 4,818 2,147 (67) 1,091,130 6,485,270 640,600 2 1,221,956 1,491 (1) 382 259 (1) 970,832 3,329,931 573,500 3 1,199,268 0 (0) 241 242 (2) 833,140 2,317,361 127,215 4 473,140 0 (0) 27 27 (4) 308,079 595,720 38,675 5 465,389 392 (149) 241 227 (74) 375,444 1,003,463 15,222 6 423,107 81 (46) 104 13 (13) 280,045 1,303,112 63,856 7 272,557 0 (0) 27 39 (39) 272,559 600,876 16,368 8 99,020 43 (11) 399 3 (2) 75,742 210,831 13,589
We also investigate the number of poisoning transfers over time. Rather than focusing on the aggregate number, we present the number of attacks from each group. Figure 4 illustrates the number of poisoning transfers (not transactions) per week from the top seven groups. We started to observe attacks from Dec. 2022, starting with Group 7. Early in 2023, Groups 3 and 4 enter the market and Group 1 comes around Feb. 2023. As we will see in Table 4, we could not confirm an early mover advantage (i.e., later comers are still profitable). The variation in the attack strategies and the active periods across attack groups further supports our clustering results.
Figure 4: Number of poisoning transfers from each attack group over time (weekly basis)
6 Analysis
We next analyze the detection outcome to identify the population(s) targeted, illuminate attacker strategies, investigate success conditions, and estimate attack profitability.
6.1 Targeted population
We first look at victim profiles, so that we can better understand who the attackers target, and eventually propose defenses. To make the attack more profitable, we conjecture that attackers are more likely to target rich addresses. As different attack groups may employ different strategies, we look at victims on a group-by-group basis on Ethereum. For this purpose, we maintain an Ethereum archive node that retains all historical account states. Specifically, we randomly sample 1,000 victims from each of the top seven attack groups. For each victim, we retrieve the sum of their three stablecoins (USDT, USDC, DAI) balance just before when the victim receives the first poisoning transfer from each attack group. We also create a baseline group by randomly sampling 1,000 USDT users and looking up their balance before the most recent transfer. Figure 5 illustrates the cumulative distribution function (CDF) of the total stablecoin balance for the victims in each of the seven attack groups and the baseline USDT users. All attack groups tend to target addresses that have significantly more stablecoins than regular USDT users. Groups 6 particularly target rich accounts given that most of the victims own more than 10,000 USD at the time of the attack. Targeting rich accounts seems successful since we detect 83 payoff transfers with more than 100,000 USD.
Figure 5: CDF of targeted victims’ stablecoin balance for the top 7 groups (and the baseline). $x$-axis is in logscale.
Attackers also appear to target addresses that are active or have transferred a large amount. To verify this, we sample three sets of 100,000 accounts: 1) the most active USDT users in terms of number of transfers, 2) the users who have made the largest USDT transfers and 3) a baseline of randomly selected USDT users. We calculate the number of attacks each user receives. Figure 6 is the CDF of the number of attacks for three groups. Active users and large transfer users appear to significantly attract more attacks than other users.
Within the targeted victim population, we confirm that the larger the number (or the amount) of transfers they are involved in, the more they are targeted. Specifically, for every victim, Spearman’s rank correlation (robust to outliers) between the number of stablecoin transfers a victim has made and the number of attacks received is $\rho=0.7$. Spearman’s rank correlation between the maximum amount of a transfer and the number of attacks received is $\rho^{\prime}=0.45$. All of those results support that attackers appear to target users with 1) high activity, 2) large transfers, and 3) high balance.
Figure 6: CDF: the number of attacks received for three groups: active, large transfers, baseline. logscale $x$-axis. The more the curve is to the right, the more the group is targeted.
6.2 Attack strategies
We next infer attack strategies. In particular, we investigate how attackers select which (and how similar) lookalike addresses to generate and target multi-chain addresses.
First, we discover that the attacker tends to imitate addresses ($\mathcal{R}$) from centralized entities. By manually looking at the list of addresses in $\mathcal{R}$ on Ethereum that the attacker imitates to generate $L$ the most, we find that those addresses often belong to centralized exchanges (e.g., Coinbase, Binance, Bybit) probably because they are one of the most active accounts. However, most addresses are hot wallets, with which normal users typically do not interact. Indeed, we do not find any successful attack involving those hot wallets. This appears to be a waste of time and money, as excluding such addresses would reduce address generation and transaction costs. We have more detailed information in Appendix E.
Second, we observe distinctive patterns in generating lookalike addresses. Figure 7 illustrates the distributions of lookalike addresses based on the number of digits matched at the beginning of the string ($x$-axis) and the end ($y$-axis) for Group 1 (left) and all other groups (right). The bubble size represents the number of lookalike addresses. The difficulty of matching addresses exponentially increases by a power of 16 for each additional hexadecimal digit (see details in §7). For most of the groups (the right figure), the bubble size decreases when $x$ and $y$ increase, forming a triangle shape starting from $(x,y)=(3,4)$ as expected. For Group 1, we observe an increase in the number of lookalike addresses starting from $(x,y)=(7,6)$ (i.e., the second triangle), indicating that Group 1 seems to use two different address generation methods. One possible conjecture is that they target different services. Most wallet software (e.g., MetaMasks, Phantom Wallet) used to show only the first and last 3–5 characters, while blockchain scan services (e.g., Etherscan) used to display the first and last 6–7 characters.
(a) Group 1
(b) Other groups
Figure 7: Number of lookalike addresses for which the first $x$- and the last $y$-digit match the intended recipient address.
Figure 8 depicts the distributions of lookalike address similarity for the top four groups. While most groups generate addresses with a maximum of 14 matching digits, Group 1 manages to produce addresses with 20 matching digits, which requires higher computational power by several orders of magnitude. The difference in distributions also suggests that different attack groups may have employed different infrastructures, which we will investigate in §7.2.
Figure 8: Distributions of the lookalike address similarity (for the top 4 attack groups): y-axis is on a log scale
Finally, we investigate whether any attack group re-uses lookalike addresses across different chains. For any EVM-compatible chains, users can use the same wallet addresses across different chains. Attackers can reduce computation especially as the most active accounts frequently hold (ERC-20 and BEP-20) tokens across chains. We observe that attackers re-use 16,903 lookalike addresses and target the same 107,542 victims across two chains. In particular, Group 4 re-use 21, Group 6 re-use 317, and Group 7 re-use 14,778 lookalike addresses between Ethereum and BSC, while the other groups do not. These results imply that the attack can easily scale up to any EVM chain. More generally, blockchain address poisoning can affect any blockchain whose addresses are represented by strings or objects complex enough that a user cannot easily spot minor differences. For example, there have been attacks on blockchains, such as Tron, [^9] https://support.token.im/hc/en-us/articles/12967949725593-Security-Alert-0-USDT-transfer-scam which do not use a hexadecimal addressing system.
6.3 When is the attack successful?
We next look into which conditions are satisfied for an attack to be successful. We test our hypothesis that the more similar the lookalike address is to the intended recipient address, the more likely the victim will make a mistake. In Figure 9, the bubble size represents the success ratio (the number of successes to the number of lookalike addresses) depending on similarity. We do not consider the number of attack attempts for each victim. While $(x,y)=(7,10)$ and $(x,y)=(9,8)$ have high success rate, there is no strong linear correlation between similarity and success ratio in general. This result suggests that the attack may have been more profitable to generate a larger number of slightly less convincing lookalike addresses and conduct more attacks.
Figure 9: Success probability and address similarity. The first $x$- and the last $y$-digit match the intended recipient address.
We next focus on successful cases and examine how address similarity (or attack timing) relates to success. Figure 10 (left) is the number of lookalike addresses targeting each victim. In 79% of the successful attacks on Ethereum, multiple lookalike addresses target one victim, indicating competition among attackers. This competition allows us to investigate how a victim selects one particular lookalike address (i.e., a winner $L_{\text{win}}$) over a pool of other lookalike addresses (i.e., victims’ perception). We focus on two aspects: how similar the winner $L_{\text{win}}$ is to $R$; and how early the winning address appears, compared to other lookalike addresses. For each successful attack, we extract the list of corresponding lookalike addresses $\mathcal{L}$, then check how $L_{\text{win}}$—the recipient of the payoff transfer—ranks in these two metrics. A smaller rank implies a higher similarity; and an earlier attack, respectively. Figure 10 shows that, while the number of competitors (left) decreases somewhat linearly, a disproportionately large proportion of the most similar addresses (center), and/or early-comers (right) tend to win the competition. In other words, in times of stronger competition, being the first and/or having a lookalike address most similar to the intended recipient seems to yield measurable benefits.
While those results provide insights from chain activities, the success also relies on what wallets or chain scanning services victims use. Appendix 8 compares how poisoning attacks appear in those services’ UI and qualitatively discusses when and where victims are most likely to fall for the attack.
Figure 10: Histogram for the number of lookalike addresses per victim (left), the ranking of the $L_{\text{win}}$ in terms of similarity (middle), and timing (right).
6.4 Benefit-cost analysis for attack groups
We next evaluate the attack profitability at the group level. Table 4 illustrates 1) the number of successful attacks, 2) the total revenue, 3) the total cost, and 4) the total profit for the top eight attack groups. The revenue is the total amount of payoff transfers (i.e., phished amount) in USD. The cost is 1) the transferred value from the tiny transfer attacks (sent from $L$ to $V$) and 2) the transaction fees for all poisoning transfers. We defer the discussion of the address generation costs to §7.2, and focus here on transaction-related costs.
To accurately reflect the price fluctuation, We use the ETH-USD conversion rates from the CoinGecko API for the day of the transaction. The profit is the difference between the revenue and the cost. To highlight attack profitability for all attack groups, Figure 11 illustrates the relationship between costs ($x$-axis) and revenues ($y$-axis) where the number is the group ID; the graph is in log-log scale. We only include attack groups with at least one payoff transfer. Revenues tend to rise with cost, which is primarily dependent on the number of poisoning transfers. The red line is the breakeven line—when costs and revenues are the same. Most groups are above the line, and thus profitable, which indicates strong incentives to carry out blockchain address poisoning attacks.
Table 4: Attack group profitability
Group Nr. Success. Revenue Total Cost Profit Attacks (USD) (USD) (USD) 1 440 12,494,248 8,368,313 4,125,935 2 363 29,043,402 2,704,455 26,338,947 3 277 9,415,636 2,546,143 6,869,493 4 109 1,187,927 504,628 683,299 5 48 4,837,606 202,868 4,634,738 6 125 9,023,888 682,016 8,341,871 7 156 4,572,499 226,418 4,346,081 8 7 103,417 68,123 35,294
Figure 11: Attack profit vs cost; red line: $y=x$
Figure 12: Win-loss ratio between attack groups
In general, the top seven groups are largely profitable. Despite the high variance in payoff amounts as shown in Table 2, those groups manage to amortize the revenue across many payoff transfers while some small groups do not. Given the attack scale (i.e., Group 1 has invested over 8 million USD), some groups appear to be well-funded cyber-crime entities. However, none of the attack instances comes from any sanctioned entities in the SDN list [41].
We also investigate competition between groups. For each successful attack, we collect the set of lookalike addresses from each group and record which group wins over the others. Figure 12 is the win-loss matrix; the color indicates the win ratio for the group on the $y$-axis against the group on the $x$-axis. The win-loss ratio of the top seven groups hovers around 50% within themselves. On the other hand, these top groups mostly win against the smaller attack groups (Groups 8–10), probably because larger groups tend to generate more (or have more similar) lookalike addresses and attack attempts.
In summary, while some attack strategies (e.g., targeting wealthy accounts, cross-chain attacks) are effective, others (e.g., imitating hot wallets, generating lookalike addresses with large $d$) are not always optimal, suggesting that the attack damage could have been even more severe.
7 Lookalike address generation
Figures 7 and 8 illustrate distinct distributions in lookalike address similarity between Group 1 and others, suggesting vast differences in computational capabilities. We next mathematically formulate the address generation problem, and simulate and measure the number of address per second (APS) of various hardware and software implementations.
7.1 Mathematical formulation
We assume that the attacker wants to match the first $a$ and the last $b$ digits of the intended recipient and maximize the total number of digits $a+b=d$, with no preference for $a$ over $b$. For instance, the attacker is indifferent between $(a,b)=(3,6)$ or $(4,5)$ as long as $(a\geq 3,b\geq 4)$ holds, per §4.1. The attacker has the set of target addresses ($\mathcal{R}$) they want to imitate.
We assume that the probability of generating a given address follows a uniform random distribution over the entire address space, and that the number of target addresses $|\mathcal{R}|$ is much smaller than the total address space ($16^{40}$). If $|\mathcal{R}|=r$, the probability of generating a target address is
$p=1-\left(\frac{16^{d}-1}{16^{d}}\right)^{r}\ ,$ (1)
which reduces to $p=1/16^{d}=2^{-4d}$ when $r=1$ (i.e., a single address is targeted).
The probability of finding a collision after $k$ trials follows the geometric distribution $P(X=k)=p\cdot(1-p)^{k-1}$, and the expected number of trials for a collision is $E[X]=1/p$. Using a first-order Taylor approximation in Eqn. (1) yields $p\approx r/16^{d}$, and the resulting expected number of trials is $E[X]\approx 2^{4d}/r$. In short, the collision probability linearly increases with the size of the target address space $r$, and, accordingly, the expected number of trials linearly decreases with $r$.
7.2 Simulating address generation
We implement and run, on various combinations of hardware and software, a brute-forcing process to simulate lookalike address generation. This allows us to measure the number of addresses per second (APS) that can be generated, which in turn gives us a lower bound on the computational resources attackers must possess to produce the lookalike addresses we observed in the wild.
We implement 1) a naive version using Python with the web3py [^10] https://github.com/ethereum/web3.py library; and 2) an optimized version leveraging CPU multiprocessing using Python with the multiprocessing, coincurve, [^11] https://github.com/ofek/coincurve, Python bindings for a heavily optimized C library for elliptic curve operations. and pycryptodome [^12] https://github.com/Legrandin/pycryptodome/, the default hashing package used by web3py. libraries. In addition, we use two popular vanity address-generating software: Vanity-ETH, [^13] https://vanity-eth.tk/ a Javascript-based web application, and Profanity2, [^14] https://github.com/1inch/profanity2 a C++-based GPU implementation. We elect, for simplicity, not to check whether the generated address is in the set of target addresses $\mathcal{R}$, which will produce an optimistic bound on performance, and thus a conservative bound on the attackers’ implied hardware capabilities.
We test our implementations on two hardware configurations, a Macbook Air (OS Version 13.1) with an 8-core Apple M1 base chip; and a server workstation (Ubuntu 22.04 LTS) with an Intel(R) Xeon(R) Silver 4214 CPU and an Nvidia RTX 8000 GPU. We run our simulations on the CPU on both machines and on the GPU on the second machine. We summarize our simulation results in Table 5.
The naive Python implementation which models what might be an expected “first-attempt” by an attacker produces in the order of $10^{3}$$10^{4}$ addresses per second. The more optimized implementation produces an order of magnitude improvement ($10^{4}$$10^{5}$). The Vanity-ETH implementation is much slower than the standalone clients. The CUDA implementation from Profanity2 is three orders of magnitude faster than our best CPU implementation with roughly half a billion APS. The Profanity2 developers self-report about a billion APS on similarly priced consumer hardware. However, the overhead imposed by checking set inclusion in $\mathcal{R}$ should be much higher in the GPU implementation.
Table 5: Generated addresses per second (APS)
Mac M1 Server (CPU-only) Server (GPU) Python (naive) 12,152 5,769 Python (optimized) 81,660 460,665 Vanity-ETH $\approx$ 4,800 Profanity2 516,437,000
Based on the distribution of the matching digits $d$ for each attack group, we next attempt to infer the attacker’s hardware capability. We define two units: CPU-day and GPU-day, the number of addresses the attacker can generate in one day per machine with a CPU and a GPU, respectively. From the above benchmarks, 1 CPU-day is $3.98\times 10^{10}$ addresses and and 1 GPU-day is $4.46\times 10^{13}$, respectively. In the process of producing a $d$-digit match, the attacker can generate many addresses matching $d^{\prime}<d$ digits in $\mathcal{R}$. Therefore, estimating the amount of computation needed to generate the maximum match $d$ for each group suffices.
As shown in Figure 8, Group 1 produces a maximum 20-digit match while other groups (Groups 2, 3, and 4) achieve 14-digit matches. We calculate the number of expected trials for $d=20$ and, based on Table 3, $r=10^{6}$. We estimate that Group 1 have used $3.0\times 10^{7}$ CPU-days or 27,093 GPU-days.
Group 1 would have needed to run over 41,000 CPUs around the clock over our whole two years of measurements to generate their lookalike addresses. We thus suspect they perform a substantial amount of their computation via GPUs. Group 1 could also have developed specialized hardware for address generation (e.g., ASICs), which would further reduce generation time. Other groups ($d=14$) only used 1.81 CPU-days, or $1.6\times 10^{-3}$ GPU-days.
To get a rough idea of the economic cost of computation, we use the AWS pricing model for similar GPUs and CPUs [^15] GPUs: https://aws.amazon.com/ec2/instance-types/g4/ and CPUs: https://aws.amazon.com/ec2/pricing/on-demand/. For GPU, an hourly rate (yearly plan) for g4dn.16xlarge (NVIDIA T4 tensor) is \$2.612 per hour; \$62.69 for one GPU-day. For CPU, an hourly rate (on-demand) for r6i.4xlarge is \$1.008; \$24.19 for one CPU-day. Using these numbers, we estimate that Group 1 would have paid (at most) 1.7M USD to generate a 20-digit match, which is well below its 4M USD profit described in Table 4. Furthermore, these numbers are likely higher than the costs incurred by self-hosting hardware. For other groups working with $d=14$, CPU costs ($1.81\times 24.19\approx 43$ USD) are dominated by transaction costs.
8 Mitigations
We next suggest a few countermeasures to mitigate the success of blockchain address poisoning.
Protocol-level mitigations: One mitigation is to map human-readable strings to each address, like DNS for IP addresses. Some entities already have adopted domain names on wallet addresses through the Ethereum Domain Name System (ENS). Human-readable addresses would reduce address poisoning as well as accidental transfers (Appendix C) and enhance usability. ENS, however, fosters other security risks such as typosquatting, name hijacking (i.e., re-using the expired domain names) [44], or name collisions [14].
Another protocol-level solution is to raise the cost of generating an address and make the attack economically less effective [29]. Adding 1 ms of latency per address results in (roughly) decreasing lookalike generation by 1,000 addresses per second. Verifiable delay functions [4] might be a good application for this purpose. Alternatively, the chain could use a larger alphabet representation to increase the cost of finding lookalikes—e.g., Bitcoin uses Base58, so the cost of matching $d$ characters is $O(58^{d})$ as opposed to $O(16^{d})$ in EVM-based chains. These additional costs could help render the attack model economically impractical.
Contract-level mitigations: One solution is to modify the major stablecoin contracts and ask permission from the token sender even when the amount is zero. Function modification requires each ERC-20 token issuer to upgrade its smart contract despite limited incentives for developers. This solution only mediates zero-value transfers but not tiny and counterfeit token transfers. In addition, token owners can quickly blacklist the attacker’s address. For the largest payoff transfer we observe, the Tether team (USDT) blocked the attacker’s address within an hour, which prevented the attacker from moving the retrieved funds to another address. [^16] Transaction blocking $L$: 0x2675aaf5db84f9f966c431ff82f2173d3251020b11b4939b4961ebedf92c78dd.
Wallet-level mitigation: Blockchain wallet or chain scanning services can improve usability. For example, one can show more of the address to make lookalike addresses easier to spot. One can also hide suspected poisoning transfers (e.g., using our detection algorithms) from the account history to prevent users from mistakenly selecting a lookalike; or at least require confirmation when sending money to a lookalike address.
User-level mitigations: Finally, users are the last line of defense. They could, for instance, be trained to engage in safer practices, e.g., building an allow-list of trusted addresses, or installing third-party extensions to flag phishing addresses [47]. This, however, adds another layer of trust, relies on the extension’s effectiveness, and remains potentially vulnerable to anti-detection strategies [23].
9 Related work
The surge in blockchain popularity exposes users to a range of cryptocurrency scams and online misconduct. Examples include phishing websites [12], giveaway scams [19], counterfeit tokens and rug pulls [8, 43, 46], market manipulation [45, 18, 11], exchange/marketplace scams [20, 35], and Ponzi schemes [39]. The above attacks exploit 1) the absence of an intermediary [22], 2) pseudonymous payments for scammers or criminals [5, 30], 3) large price movements attracting inexperienced investors [31, 15] and inviting mischief [36, 17].
We highlight work particularly related to our discussion, with a focus on attack scale. Gao et al. [8] identified 2,117 counterfeit tokens by applying keyword matching to the 100 most popular tokens. The authors investigated two techniques involving counterfeit tokens (airdrop and arbitrage), leading to total financial losses exceeding 17M USD over 7,000 victims. Xia et al. [43] identified around 10,000 counterfeit tokens on Uniswap (50% of all Uniswap tokens) through ground truth labels provided by the platform, address clustering, and machine learning-based techniques. They estimated these counterfeit tokens collectively led to financial losses of 16M USD from approximately 40,000 victims. He et al. [12] introduced TxPhishScope, a tool designed to detect blockchain phishing websites. The authors identified 26,333 phishing websites with 3,486 phishing blockchain accounts. Other studies on free giveaway scams reported 24M–69.9M USD losses [19], including 872K USD losses from Twitter [17]. Based on these other works, blockchain address poisoning appears to be one of the largest (in terms of aggregate losses) and most wildly targeted criminal activity taking place on-chain.
Closest to our efforts, Ye et al. [46] examined three types of blockchain scams, one of which is address poisoning on Ethereum with zero-value transfers. In parallel, and contemporary to our work, Guan and Li [10] also studied address poisoning on Ethereum and discuss all three poisoning transfers. We provide a detailed comparison with our method and results in §4.1, §5.2 and Appendix A. The Chainalysis’s blog post [34] conducted the case study (WBTC 68M loss) and analyzed its related addresses (victims, lookalike addresses).
10 Conclusion
We studied blockchain address poisoning, a form of phishing that exploits visual similarities in hexadecimal wallet addresses used in EVM-compatible chains (e.g., Ethereum, BSC). We implemented a detection, identified large attack groups and their strategies, and inferred their capability through simulation. We showed blockchain address poisoning presents a considerable threat, in total, 270M attack attempts, 50M lookalike (malicious) addresses, and, more chillingly, nearly 84M USD in monetary losses (from more than 6,600 cases) on Ethereum and BSC. Worse, damages could be even higher if attackers optimized their strategies.
11 Ethical considerations
The concept of blockchain address poisoning is already the subject of some online discussions, and, as this paper shows, this issue is currently extensively exploited in the field. As such, wallet operators and (some) users are most likely already aware of at least the theoretical possibility of the attack. What our work does, though, is to demonstrate its practical severity in economic terms – and we hope it will be a call to arms to act sooner rather than later.
While attackers could potentially misuse some of our findings to improve their strategies, besides raising community awareness, the main contribution of our paper is to extensively describe the attack mechanisms. We also introduce possible protection strategies. Our proposed mitigations can hopefully inspire wallet operators. Likewise, we hope that our work can help users recognize when they are in targeted populations, and accordingly, take precautionary measures.
Finally, Appendix G, specifically Table 9, includes full transaction hashes, which could potentially be used to re-identify the wallet addresses of the potential victims (whom we do not directly list in the table). We decided to present the table as is, as obfuscating part of the transaction hash would impede reproducibility or external validation of our findings, for little to no privacy benefit: users of Ethereum (or BSC) are relying on a public blockchain, in which all records are publicly available. Furthermore, wallet addresses are not, on their own, considered personally identificable information, as 1) there is no one-to-one mapping to specific individuals, as a given person may hold several wallets, and a single wallet may be shared by multiple people (e.g., belonging to the same organization), and 2) a wallet address does not, in itself, contain information directly mapping to a real-world identity.
Acknowledgments
This research was partially supported by a Sui Foundation Academic Grant, the Carnegie Mellon CyLab’s Secure Blockchain Initiative, and the Nakajima Foundation.
Appendix A Comparison with Guan and Li’s
This section compares our detection performance to the most recent work from Guan and Li [10]. To align with their observation, we only include our measurement results from Nov. 2022 to Feb. 2024 on Ethereum. Table 6 summarizes the results.
First, the total amount of payoff transfers is similar, corroborating the validity of both approaches. Second, both approaches estimate similar attack costs in ETH, but not in USD. Guan and Li use a fixed ETH to USD conversion rate (around 3200 USD) whereas we query the price at the time of attack. The exchange rate, however, fluctuated between 1000 USD and 2000 USD for the majority of the observation period, leading to the attack costs being overestimated by over 10 million USD, which results in a difference of a 300% on the return-on-investment (RoI) for the attacker. Third, we detect more than 2.2 million counterfeit token transfers. Guan and Li do not report tokens with symbols that do not exactly match USDT and USDC. For example, out of 5,838 counterfeit tokens, only 1,922 use the exact same token symbols whereas at least 740 tokens use the different encodings of USDT or USDC (e.g., using a Cyrillic “S” and/or “T” in “USDT”). Likewise, attackers can also avoid their detection if they change the attack characteristics, e.g., launching zero-value transfers without attack contracts. Our results suggest filtering the transfers based on address similarity first would be more robust to attacker manipulation (less false negatives), than filtering by attack characteristics. On the other hand, compared to Guan and Li [10], our technique misses attacks where less than 7 digits match the address $L$.
In addition to the methodological differences, our work has several distinctions. First, we have better coverage in measurement (two years, two chains). We find 270M attack attempts (compared to 14M), highlighting significant chain waste and user experience degradation.
Table 6: Comparison of detection performance
Payoff Cost ETH ROI Tiny Zero-value Counterfeit CT transfer (USD) token Guan and Li [10] 76.79M 7918.2 (25.5M) 2.2 277.3K 7.85M 6.64M 2,333 Ours 76.81M 8202 (14.9M) 5.1 283.9K 7.05M 8.87M 5,838
Second, although Guan and Li also perform attacker clustering, they appear to have overestimated group sizes, due to erroneous merges from copying bots, as we show in §5.3. We carefully remove bots and extrapolate a few large attack entities.
Last but not least, most of our analyses are novel. We analyze victim characteristics, success conditions, group competitions, cross-chain attacks, and modeling of address generation for the first time. We provide deeper insights on how attackers perform attacks and what capabilities they have.
Appendix B Birthday Paradox
This section investigates whether our detection system mislabels poisoning transfers when the victim interacts with two similar-looking addresses by coincidence (i.e., due to the “birthday paradox”). We define the number of accounts the victim interacts with as $r=|\mathcal{R}|$ and the likelihood of encountering, among these accounts, by pure chance, two addresses with the first three and the last four characters matching as $p$, a “collision probability.” By the birthday paradox, we can mathematically relate those two variables with: $p\approx 1-e^{{-r^{2}}/{2\times 16^{(3+4)}}}$. For instance, when the victim interacts with $r=$19,290 accounts, the probability $p$ becomes 0.5.
Here, we examine whether our system should exclude a set of victims that present a collision probability greater than a given threshold $a$, i.e., $p\geq a$. We denote this set of victims by $\mathcal{V}_{p\geq a}$. Figure 13(a) shows the number of victims ($|\mathcal{V}_{p>a}|$) that present a collision probability greater than $a$, on the $x$ axis. Even at a small threshold $a=0.01$, we only find 664 accounts (0.04% of total victims). At first, excluding $\mathcal{V}_{p\geq a}$ does not seem to impact our detection performance. However, we find that excluding these accounts results in numerous false negatives because attackers purposefully target very active accounts, as shown in §6.1. Figure 13(b) illustrates the number of attacks we would report if we excluded attacks on $\mathcal{V}_{p>a}$. For example, we would discover 16.2M million poisoning transfers at $a=0.01$, which is 1 million fewer than at $a=0.99$. Indeed, we find that most of those detected attacks are indeed poisoning transfers. 99.7% of attacks targeting $\mathcal{V}_{p>0.99}$ are launched by the attack accounts $\mathcal{A}$ that have also targeted victims $\mathcal{V}_{p<0.01}$ (i.e., the population that is hardly susceptible to a collision).
False positives on payoff transfers are even more unlikely as they require all of the following: 1) the victim interacts with two similar addresses, 2) one of similar addresses “unintentionally” performs the poisoning transfer (e.g., testing with a zero transfer), and 3) the intended transfer, the poisoning transfer, and the payoff transfer happen in that correct order. As shown in §5.2, we do not find any false positives for top 30 cases for both chains.
The result shows that our system is robust against spurious collisions and that setting a low threshold reduces the performance of our detection system. Hence, we decide to set the threshold of $a=0.999$ to exclude victims that are interacting with a large enough number of addresses to be extremely susceptible to collisions, while capturing nearly all attacks. Setting such a high threshold only excludes, for instance, 64 Ethereum accounts.
(a) Nr. of accounts ($y$-axis) with a probability of collision greater than a given threshold ($x$-axis)
(b) Nr. of poisoning transfers detected at a given collision probability threshold
Figure 13: The number of victims that are susceptible to a collision and the number of attacks that are detected when we exclude those victims. The $x$-axis does not start with 0.
Appendix C Accidental transfers
Victims could accidentally send tokens to an incorrect, but not malicious, destination address, due to a typing mistake. We focus on those cases when two addresses (legitimate and lookalike) match more than 20 (out of 40) characters, regardless of position. Given current hardware capabilities, generating a lookalike address with more than 20 matching digits is computationally unlikely (see §7 for details). Furthermore, if users mistakenly specify an incorrect destination address, tokens are probably not recoverable, since it is very unlikely that anybody possesses the private key of the mistyped address. We leverage this property, and check whether one of the recipient addresses has ever sent any authentic tokens; [^17] To be more precise, we exclude zero-value transfers or counterfeit token transfers which can be launched without private keys. if they have not, we hypothesize this transfer is an accidental transfer that we categorize separately from payoff transfers.
Measurement results. We identify 363 accidental transfers (i.e., not the result of a deliberate attack) from 295 victims and 336 mistyped destination addresses, totaling a 5.5M USD loss in Ethereum. This high number suggests that some victims have sent assets to the same mistyped address multiple times, or made typing mistakes multiple times, over different intended addresses. We observe a similar number of accidental transfers in BSC: 318 cases from 206 victims and 241 mistyped addresses for a total of 57,562 USD lost. By manually looking at accidental transfers, we confirmed that those transfers appear to be typing mistakes, and noticed common human errors. For instance, users 1) mistype a character with a neighboring key (i.e., low fat finger-distance [21]): “4” vs. “3,” “3” vs. “e,” 2) swap the order of two characters, or 3) mistype a character for something visually similar [32]: “b” vs “d,” “e” vs “c,” or “6” and “b.” To quantitatively assess typing mistakes, we calculated the Levenshtein distance $\ell$ (i.e., the number of operations necessary to replace a string with one another, including transposition) between two addresses. As expected, the distance is small; 84.95% of typos have $\ell=1$, and 96.69% have $\ell\leq 2$ in Ethereum. We have similar results in BSC (86.48% and 94.34% for $\ell=1$ and $\ell\leq 2$).
Appendix D Detecting copying bots
Here we explain why we use the attack ratio to exclude bots and show the robustness of our clustering results through temporal clustering.
As discussed in §4.2, we group two poisoning transfers if they 1) are in the same $TX$, 2) are launched by the same attacker $A$, or 3) specify the same lookalike address $L$. Our cluster considers two groups to be the same if they share even one link. We notice that two seemingly distinct large groups (based on their strategies and behaviors) are connected by a few attack addresses $A$ that behave differently from other attack addresses. We find that they appear to be the copying bots that duplicate transactions from multiple attack groups. For instance, the account 0x7D575a7C732D1c502c07f18BB822D29CF7DBf9E8 copies not only address poisoning transactions [^18] The original transaction: 0x725eaedf8857e243587020de97e-
d503fb8bd8899bcbe8c685cf57fbce6810cc5
the copy transaction: 0x5cc772d10b9e6fd61c7402967b5ae2ff6f-
cae7a2ca9bef0e9c1c6202c2ea257c
but also other transactions such as from MEV (Maximum Extractable Value) bots [^19] The original transaction: 0x0720fec6980c196b3545d273da2-
b5b41987bf29524c6236cbf462d683df06057
the copy transaction: 0x7aec84ca51fd5bdfcecf8717762-
87b6732e8b8f27097e8d690c090cf6e6865bf
. This type of copying behavior is relatively common in Ethereum or BSC, as discussed in previous work [27, 26].
Those bots seem to copy the original transactions without properly understanding address poisoning. Without changing the lookalike address $L$, the payoff transfer goes to the original attacker regardless of who initiates the attack. In other words, these bots are helping other attackers.
We next explain why we choose to use the attack ratio to detect copying bots over other methods (using the notion of time or removing all duplicated transactions).
Using the notion of time. Copy transactions generally come after the original poisoning transactions (in terms of the block number), so removing the second or later transactions could be effective. However, we notice that some copy transactions end up in the same block as the original transaction. This can be achieved by monitoring unconfirmed transactions in the peer-to-peer (P2P) network. For instance, the account 0xfffdcF2B3419C243A5eba5f051A64ad629362c9a manages to include copy transactions in the same block as the original transaction most of the time. Based on the public mempool data from Flashbots [^20] https://mempool-dumpster.flashbots.net/index.html, we realize that this account’s transactions do not appear in the public P2P network, while the original transactions do. This suggests that some bots monitor the original transactions in the P2P layer, and directly submit their transactions to the block builders (to potentially front-run the original transaction). As a result, we cannot systematically determine whether the first transaction is the original or the copy.
Removing all duplicated transactions or transfers. While differentiating the original from the copy transactions remains impractical, we can simply remove all transactions involved in copying behavior (including the original). If the bot simply copies the entire transaction, the data field in the transaction stays the same. While this method captures most of the copy transactions, we observe that some bots appear to slightly modify one of the parameters to potentially avoid detection, yielding a different data field.
We can also exclude duplicated transfers that have the same ($L$, $V$, $CT$ (or $AC$), and $v$ (value)). This method appears to capture all the transactions from the bots. However, we realize that the attackers (not the copying bots) also re-send their poisoning transfers several times against the same victim. We would remove more than 11 million (62%) “duplicated” transfers with many false positives, which significantly undermines the final clustering results (i.e., the size of the clusters becomes exceedingly small).
Removing potential bots by attack ratio. Only using information from transactions either leads to 1) the removal of too many non-bot transactions or 2) the inclusion of copying bot transactions. We decide to retrieve additional information of attack accounts $A$, and exclude $A$ with a low attack ratio—the ratio of the transactions used for address poisoning, as introduced in §4.2.
We aim to set the threshold as low as possible while removing all the copying bots. Figure 14 (left) illustrates the top 10 clusters when we vary the attack ratio threshold from 0 (removing no transfer) to 1 (removing all transfers). The line indicates the same cluster. We manually identify that all the erroneous merges we have spotted disappear with a threshold of 0.5; the largest cluster at threshold 0 breaks down into five different groups, which, as we show in §5.3, exhibit distinct behaviors. Figure 14 (right) shows the ratio of removed transfers over the total number of transfers ($y$-axis) for each threshold value ($x$-axis). Even setting a high threshold does not exclude many attacks, because attackers generally use $A$ only for address poisoning. The attack addresses $\mathcal{A}$ with a threshold below 0.5 account for only 1.45% of all transfers (represented by the vertical line on Figure 14 (right)), which helps avoid removing too many attack instances and preserve the structure of the clusters. Groups remain stable up to a threshold of 0.7, but the group structure starts to collapse with a threshold above 0.8 (i.e., removing 6.64% of all transfers). Therefore, we choose a threshold of 0.5.
Figure 14: Left: Top 10 clusters ($y$-axis) based on the threshold ($x$-axis). Right: the ratio of removed transfers over the total number of transfers based on the threshold
Finally, to additionally validate the robustness of our clustering, we perform a temporal clustering. We apply our clustering algorithm for the data until time $x$ and gradually increase $x$ over time to observe the change in the clustering results. Figure 15 shows how the top 10 clusters (based on the number of lookalike addresses) evolve over time where the link indicates the same attack group. The left figure shows the result with a threshold of 0.5; the clusters appear to be stable over time. The right figure is the result without any threshold (no bot exclusion); some large clusters remain separate for a while, but eventually connected around July and August 2023, forming a single large cluster due to the appearance of copying bots.
(a) With a threshold of 0.5
(b) Without a threshold
Figure 15: Temporal clustering results. The x-axis is the time, and the y-axis is the number of clusters.
Appendix E Intended addresses
We focus here on the intended addresses $R$—the victim’s recipient which attackers try to impersonate. Table 7 lists the top ten intended addresses the attackers impersonate the most (based on the number of lookalike addresses $L$ targeting each $R$). We label each address based on the third-party Arkham Intel website. [^21] https://intel.arkm.com/ We also aggregate the number of attacks targeting each $R$. The table shows that many of $R$ are centralized exchange hot addresses. Since users do not directly interact with those hot wallets, we do not observe any successful cases for those addresses. Yet, those exchanges’ deposit addresses frequently interact with those hot wallets, which may explain why the attackers’ (algorithm) capture those intended addresses (as discussed in §6.2).
Table 7: Top 10 intended addresses
Intended address $R$ Label Num of $L$ Num of $TR$ 0xA9D1e0…1d3E43 Coinbase Hot Wallet 143,764 18,229 0x28C6c0…f21d60 Binance Hot Wallet 26,373 16,147 0x974CaA…ECC400 Stake.com Hot Wallet 3,254 2,355 0xF7C8dA…38d921 Devout Deposit 5,242 2,014 0xf89d7b…5EaA40 Bybit Hot Wallet 4,556 1,702 0xBFCd86…23d6aD HitBTC Hot Wallet 3,216 1,569 0x731309…AB4D2a BitGet Hot Wallet 7,547 1,477 0x99870D…9AF1e2 HitBTC Deposit 9,700 1,449 0x3CC936…aeCF18 MEXC Hot Wallet 6,603 1,397 0x75e89d…1dcB88 MEXC Hot Wallet 1,629 1,257
Appendix F UI Design of wallets and chain scanners
This section examines how address poisoning attacks appear in victims’ wallets or chain scanners. In particular, we visit those services in Jan. 2025 and investigate how those services have reacted to the threat of address poisoning.
To discover how victims see the attacks on their wallets and chain scanners, we first implement address poisoning transfers and attack our own accounts (i.e., attacking our accounts) on the Ethereum mainnet. While our attack uses minimal blockchain resources (three poisoning transfers), conducting our attack on the mainnet is critical because of the significant UI difference from the testnet. Specifically, we conduct our experiment on Dec. 20th, 2024, and used 159,364 gas for three poisoning transfers which consumes about 1% of one block (0.00014% of gas usage for a day).
We generate a lookalike address $L$ that matches the first six and the last 5 characters to $R$. We follow to $AC$ and $CT$ implementations from the attackers who have uploaded their contracts’ source codes on Etherescan and perform three poisoning transfers on our victim account $V$. Our performed attack can be found in $V$ 0xB6e84DF1cE401117C450221ccc6EF502cb0e2284.
First, we check the victim’s UI on blockchain wallets (as of Jan. 2nd, 2025). Based on the wallet market share [25], we study MetaMask, Trust Wallet, and Phantom. Regarding the address abbreviations (first, last), Metamask is (4,4), Trust Wallet is (2, 4), and Phantom is (4, 4). Trust Wallet and Phantom only display tiny transfers but not zero-value or counterfeit token transfers. Metamask does not appear to show any ERC-20 token transfer history in the first place. When initiating the transfers, users no longer have options to select the recipient from “recently used addresses” but manually specify it on their own. The fact suggests that victims may rely more on blockchain scanning services to find the recipients.
Second, we check the major blockchain scanning services: Etherscan, Ethplorer, OkLink, and Arkm (as of Jan. 2nd, 2024). Table 8 compares the UI of those blockchain scanners (tiny: T, zero-value: Z, and counterfeit token: C). The level of address abbreviations significantly differs by each service. Etherscan equally shortens the beginning and the end while others significantly abbreviate one side (or no abbreviation at all). Since the attackers mostly focus on the first and last characters now, displaying one side might be effective. Next, we check the poisoning transfers. By default, most services (but Ethplorer) successfully hide zero-value transfers while they fail to detect tiny transfers. By changing the setting, Etherscan and OkLink additionally show zero-value, and Arkm displays counterfeit token transfers. It is generally easier for them to hide tiny than zero-value (i.e., just detecting the value of 0) while they must implement some heuristics to detect tiny or counterfeit token transfers [^22] Etherscan successfully detects all poisoning transfers as of our visit on Jan. 13th, 2025.. Note that hiding zero-value transfers hinders the usability if users use zero-value transfers for (benign) testing purposes. Most services show token icons or give a warning to $CT$ (based on their list of $AC$), which would be helpful for users to distinguish them.
Table 8: UI comparison of blockchain scanners
Addr. abbr. Poisoning transfers Token icon Default Setting change Etherscan 8, 9 T C Z Yes Ethplorer No abbr. T, Z, C No OKLink 4, 11 T, C Z Yes Arkm 36, 0 T C Yes
Appendix G Payoff transfers
Table 9 presents the 30 largest payoff transfers on the Ethereum blockchain and BSC as examples for readers to investigate. These transactions are the final step of the attack, in which victims mistakenly send funds to the attackers. We verify the authenticity of these attacks discovered by our algorithm on Etherscan (BSCscan) through the following steps:
  1. 1.
    In the transaction page, get the sender (i.e., from) of the Transfer event
  2. 2.
    Go to the sender (victim)’s page and examine the Token Transfer (ERC-20) tab.
  3. 3.
    Follow the prompt and change the site settings to disable “Ignore Tokens with Poor Reputation” and “Zero-Value Token Transfers” options.
  4. 4.
    Find the payoff transfer transaction and verify that there are attack attempts in earlier blocks. For example, in the first case of Ethereum, a transfer of 10,000,000 counterfeit tokens from the victim to the (attacker’s) lookalike address took place in block 17,778,047.
  5. 5.
    Find the authentic token transfer events from the victim to the intended address that share similar characters to the lookalike address. For example, in the first case of Ethereum, a transfer of 10,000,000 USDT (Tether) is sent from the victim to the address 0xa7B4BAC8f0f9692e56750aEFB5f-
    6cB5516E90570
Table 9: Largest payoff transfers on Ethereum and Binance Smart Chain. USD Phished is the monetary loss in this single transaction, Recipient is the attacker’s lookalike address, Intended recipient is the address that the victim interacted with, Matched is the number of characters matched between the lookalike and the intended recipient, $a+b$ indicates that the lookalike address matched the first $a$ characters and the last $b$ characters of the intended address.
(a) Ethereum blockchain
USD Phished Transaction Hash Recipient (attacker) Intended recipient Matched
20,000,000 0x08255ca0e42a872559437141fa46980e66d907f7668922467d67515b1ebb4b7f 0xa7Bf48…E90570 0xa7B4BA…E90570 3 + 6
3,800,000 0x2dea87ff34b0d4712d59e582e810d1a8396e469ce7830a61b9521759b28221f1 0x61b697…074c5C 0x61b9d3…174c5c 3 + 5
3,554,610 0x11ac26acce0620c50731a94b528419771211779903d6917b05f2bd90e4a6476c 0x1cbB23…9B758a 0x1CbB23…9B758a 7 + 6
2,030,000 0x0911bd6713493e9ab75ef82cc909114218996f0e717b81b71d7bf29fa06e1622 0x74C9bd…60E1cA 0x74C32c…50E1cA 3 + 5
2,000,000 0x48e591d562a5098527c0de850ba44ce21014726434184f693707f62877a16994 0xbb2EDb…619455 0xBB217c…219455 3 + 5
2,000,000 0x5ae3147442cb9cba98b8b3dc1548a7c27dab29847ca2c126059ca806ad4479c9 0xC7B14b…B33A8f 0xc7b57d…533a8F 3 + 5
1,999,000 0x7cae29b7d215ef67fa183d4455760fc6e8698bfb2c5f1744dec4749d0c4b88c7 0xC7B14b…B33A8f 0xc7b57d…533a8F 3 + 5
1,660,000 0x86d39a9fd0e4078223179e6f59d2651963edb07899efc9558c43a0a338fb40a5 0xEA0d13…CBeb14 0xEa0ddb…FcEb14 4 + 4
1,499,990 0xf9948a832bfc217451f056fb3dc2df258c6ad525d7c3c184e07eeeee3df7ed09 0x9339e7…1F4126 0x9339CF…3A4126 4 + 4
1,200,000 0x371f26eb4b08b37d69b280fa5ea06208750c4bc9cf3258950d63c842f9f01154 0xCbA796…34c994 0xCbA796…34c994 8 + 6
1,045,150 0xa7be1b408b17853c3f9de8cef05371616d154b646e0f7c556feb72dc87ca9676 0x73435A…ca79F7 0x734659…Ca79F7 3 + 6
1,000,000 0xb97df9feb55c3849da4551f43995ec0f0fd78ff0e3d28f337934dbaba154bc11 0x80D707…4dbeA2 0x80D713…f4beA2 4 + 4
960,000 0x5b03c7356fe51575bd4fd190cb83841ec7398d8531dd34d21d89fdc0f776afd5 0x946C8e…f98Cd6 0x9462B5…f98CD6 3 + 6
865,884 0x71eb719e26a31dbb9b05d38b8a36d73537af35e3a5e0d2ccf4bbb6accbd8cd36 0x55B714…Ba9d79 0x55be3c…Ca9d79 3 + 5
849,996 0xb09f4c9fe09e6bbeb3abfdbed21a0403a2d8d2b6313ed1cbcdbd620a1f748fb2 0xEB4034…84683d 0xEb4034…84683D 8 + 7
800,158 0xaeb6f1b730e65ddb05c84d71af8c370dead57c4d1aecafc7e2369fde50875ae6 0x0Cee6B…e70EbF 0x0CEE33…3f0EbF 4 + 4
749,900 0xecd48ffd8d6a09f1ca18e7acbaad4ff4f70116d70fa65dab842a7cd1651df8a0 0xd7F373…7B7403 0xd7F38D…7B7403 4 + 6
719,000 0x1576beff54a53f0ef0d881a6242edf4539d0964b5b2431f9e1f08fe781b42de3 0x9887E5…D40ea3 0x9887F3…740ea3 4 + 5
710,000 0x66096743f07f2f0d49818f5e4de28b98bbf3ac6bfcf665fa48225f97260be6b1 0x949D0D…ef861B 0x949954…58861B 3 + 4
699,990 0xb2d6e74cba922beeae6dfd9f8a832372ae1d4d5314c0d1505dda13911e840316 0xC7BA77…26b438 0xc7BD59…F5B438 3 + 4
618,000 0x273166dfae3d85e8297a0c1353d8d9bc72ab7b021a61797cd30b8d1871f59fea 0x55Eb2A…A83159 0x55e5Cf…083159 3 + 5
600,000 0xf997e3b1fcba63c52815dbc63d4364691a79ef65bb578957f182f54753ef59ca 0xE07222…6adFeA 0xE07C8D…6AdFea 3 + 6
561,035 0x65f005042adff7a66fd5e317459987367dcfbae3246a50dbde88d1f364c4d9a5 0x368595…c3aFd3 0x3683a1…C3aFD3 3 + 7
541,900 0x82be73572215148ba7c03e984b1361188b098b7ceedf1aaac2a2574ec3f52fa0 0x35D6CD…4B0305 0x35D3c7…4b0305 3 + 6
541,000 0x7b5b69fad190d0e2b233d1b5bc0fb4a92178b72a737b3f2041a438a8bcd8ecad 0x60855B…30D466 0x6085fd…c6D466 4 + 4
500,962 0xea71be058fcd162620d95d04b2d5acbec953cedbb5ccb5339e6d43198b4b026e 0x539592…0DDf0b 0x539683…a6Df0B 3 + 4
500,000 0x8e452c91d99185a0ba325b2ea3dbb276d09f4b2a6dc6772d615acd48237a3148 0x324AFa…80C28F 0x32494C…C0c28F 3 + 5
500,000 0xd4bc53f71576fc1b78dd56145e42a127d4109567e7d04c3033e59d9ce9da502c 0xB08cfc…A74a2F 0xb08D97…574A2f 3 + 5
500,000 0xe68134277191eae950ca698b6651ac7c1942c6c4578eb53a98379c7deceb3f33 0x140588…b0Cb01 0x140b09…25cB01 3 + 4
500,000 0x78cfaa6c26a06444561397b7b02a426b1e3250b1eaf7e051e91b9137a42fdf86 0xC6BCCb…Bf0115 0xc6B7A8…BF0115 3 + 6
(b) Binance Smart Chain
USD Phished Transaction Hash Recipient (attacker) Intended recipient Matched
279,489 0xa0880a04b9f4dd114eb6e5b457a4b6f260c1e1444eef350b62cdc5f1cba8b0ca 0x7e4d7a…a7f223 0x7e4D54…dcf223 4 + 4
150,000 0xba144f4c4542434bc1ebc2d70030a89f46e81c993842bf79b43ad453cd446d3e 0x112C71…C4B08a 0x112784…24B08a 3 + 5
144,876 0xc11c5930465a2c33a479b7df63eaa15610dd9ca78c70f788d0f69cfe5e96960d 0x25f566…b02c50 0x25f5e8…b02c50 4 + 6
133,000 0x266647bc5190cd516f791399ac61b10e876ad75dcc0c6fabb77ccc80f7ead038 0xe72a80…49dc98 0xe72dee…f7dc98 3 + 4
100,000 0x62798e183339f5ed70239a84b4463b7e48f497e1ef0e23ee82b78087195462d4 0xD643f1…eAF6fB 0xd64329…24F6fB 4 + 4
97,250 0x9a4081519e99e2e6fe6add185da7333d44eb28cd1c70df36a32ff8648940cbf5 0x8AC739…5cEB7B 0x8ac7CE…76eb7b 4 + 4
91,680 0x12505feedc126def4e118a2b8e970766795d45201a79dbe45f4b2b9a41391375 0x5ebefa…8De436 0x5EBe2a…EDe436 4 + 5
86,180 0xa070b46843b57ff4e529cbe7aabb67c99cb36da58510106bb464d1f039e7b3db 0x2556be…757cF1 0x255EC6…A57cF1 3 + 5
72,671 0x4e7dd743e32d51f6f8d0705d20542a15a9a8b40a80e707ad1e1f18147033b81b 0x6c9351…D01114 0x6c93Ce…EC1114 4 + 4
70,006 0x13aecc1f514cc95db38a31ada1c9827124aa87360dc11cb1b3dfd7fe0d2499e0 0x2f9233…2BC04D 0x2f9276…cbc04d 4 + 5
60,000 0x521bb785beaac9b08cc35d57096a40d28dc7e3e697fe22572a0431681770cc04 0x1616bB…4d1800 0x161c2C…6d1800 3 + 5
50,111 0xe5576fff71817e990959b6947aa54866888953e2912d3047620b27d0ac766350 0xf06e0a…b8b610 0xf06e59…29b610 4 + 4
50,094 0x09c21db09c5e92e68bf01eb46374796a9a3c16a0b42df9ef2217020f2f8e2352 0x5c3415…82E06f 0x5c3464…a2E06F 4 + 5
50,000 0x6108401848776037926010c09b848f00ec5e04e3c4075b72437575544e4c0cd9 0x809799…4A2d4c 0x8097dc…ba2D4C 4 + 5
50,000 0xb4a277c05232537bbb1dd2aa1f04c8dced24fbd5be125c65fe0498db3f4b837d 0x1487c9…f08B32 0x148799…b08b32 4 + 5
50,000 0xedfe46269453c28ff4d70f736ca73664ca032bbc65a67462b3d35ea178e47171 0xC4d43e…71d48D 0xc4D89C…81D48D 3 + 5
46,000 0xe4335fecccba5eec21dbaa18d71cdb4d1c56056ee38c8a30b2a57aaccf4c4faa 0xcc3af0…97298f 0xcc3afe…fe298f 5 + 4
40,008 0x230ef03a64905d8e0c9a074fde33c0fb993ad536b3852129c36efe485409ad69 0x458728…dF9961 0x4587C1…629961 4 + 4
39,995 0x5ac785bf1bc5273ca3183471bafce3e1a301afc969f3e5eb26e83f249cca70b4 0x73A31C…E5A703 0x73af35…35a703 3 + 5
37,000 0x41c732fa968e3816c69ffaebc849233b84a863695344f24328891654b1230edf 0x52C990…0c1961 0x52c908…fC1961 4 + 5
35,000 0xd96aa7bfbd32720c88ecfddaec21af640a65cb858ddb735d23b5d6b1fae2b1bc 0xe6287d…08d0EA 0xe62d77…f8D0ea 3 + 5
32,000 0x7cee39ef7f56a08b88346938540be29d58249a33b1b53a1166203d7c467e21c4 0xF632D3…294ee0 0xf63078…194EE0 3 + 5
30,060 0x53eb5af3962cf511e51b6369071494118460cc11e24234a3551478f3415e0a47 0x4245B2…782880 0x424276…782880 3 + 6
30,000 0x6bc9148d278302517d49f7602f5c8fd0ef516ea8563996910db846ab029e6994 0x8b8ea3…7A8F9F 0x8B808A…5a8f9f 3 + 5
30,000 0xa9c015c3b41dd449266cb67fb8c514aadc501fac88a1c49e06d6b51b0c0a9953 0x532D78…C1540f 0x532Dc7…71540f 4 + 5
30,000 0x6462aaf80cc01700133399d6086cc0381f72ae48b52cdd85529fd7ec45d19fc9 0x2FD1D3…5cc8B8 0x2fd1Da…cCC8B8 5 + 5
30,000 0x6ef40292302c45c1a0a1756b9b431e4b08f02118ca78c12892377c52e5a84796 0xBB3e51…c3aEAF 0xBb3fd3…93aEAF 3 + 5
29,377 0x74123f8aef47435e94a4c2291df65bdcf1d6735de97e05fb84974c48a207983d 0xE15D59…5bA949 0xE15f47…6ba949 3 + 5
28,650 0x1e56b900524521ef21e82392ff8992675a58f0af74a8ddf476c96876533cd5c2 0xaF64F8…Cf01d0 0xaF6859…3f01d0 3 + 5
25,666 0x60b3ca5895303d08fe83288419e805de7cd078b1778ba7ee30ceafdbac170601 0x4A2274…7b57A0 0x4a22aE…bB57a0 4 + 5

### v3

@@
 Compared to contemporary studies, this paper provides four notable contributions.
-First, we develop a detection system and perform measurements over two years on Ethereum and BSC.
-We identify 13 times the number of attack attempts reported previously—totaling 270M on-chain attacks targeting 17M victims.
+First, we develop a detection system and perform measurements over two years on both Ethereum and BSC.
+We identify 13 times more attack attempts than reported previously—totaling 270M on-chain attacks targeting 17M victims.
 6,633 incidents have caused at least 83.8M USD in losses, which makes blockchain address poisoning one of the largest cryptocurrency phishing schemes observed in the wild.
@@
 Third, we reveal attack strategies—targeted populations, success conditions (address similarity, timing), and cross-chain attacks.
-Fourth, we mathematically define and simulate the lookalike address-generation process across various software- and hardware-based implementations, and identify a large-scale attacker group that appears to use GPUs.
+Fourth, we mathematically define and simulate the lookalike address generation process across various software- and hardware-based implementations, and identify a large-scale attacker group that appears to use GPUs.
 We also discuss defensive countermeasures.
@@
 (c) Counterfeit token transfer
-Figure 1: Poisoning transfers
+Figure 1: Three types of poisoning transfers
 Most modern blockchains rely on “wallets” for monetary transfers.
@@
 We design a detection system and run it between July 1, 2022, and June 30, 2024, on Ethereum and Binance Smart Chain (BSC).
-Compared to contemporary studies [46, 10], our detection 1) covers all types of poisoning attacks, 2) demonstrates greater robustness against attacker manipulation, and 3) generalizes better across time intervals and chains.
-Combining the results from both chains, we identify over 270M attack attempts (i.e., 13 times the previous efforts: 21M [46] and 14M [10]) that target over 17M victims.
+Combining both chains, we identify over 270M attack attempts (i.e., 13 times more than the previous efforts: 21M [47] and 14M [10]) that target over 17M victims.
 Out of those, attackers successfully received 6,633 transfers, adding up to over 83.8M USD in ill-gotten gains.
 This attack appears to be one of the largest cryptocurrency phishing schemes observed in the wild.
-We validate our results by 1) manually verifying the phished cases, 2) directly comparing the results with the related works, 3) checking the smart contracts used for those attacks, and 4) comparing the data collection results with different parameters (e.g., a window size).
-Of independent interest, we leverage our detection system to capture accidental transfers (e.g., typing mistakes in the recipient’s address).
-We confirm 681 accidental transfers with 5.5M USD losses, based solely on human mistakes which underscores the need for more forgiving user interfaces.
-Attackers typically optimize their operations by bundling several transfers into a single transaction and reusing both addresses and contracts.
-This behavior allows us to link attack addresses through the “guilt-by-association” heuristic [43].
+We evaluate our detector by 1) the external dataset (100% precision, 97.2% recall), 2) analyzing its false negatives, 3) manually verifying the phished cases, 4) checking the smart contracts used for those attacks, and 5) comparing detection outcomes with different parameters (e.g., window size and similarity threshold).
+Attackers typically optimize their operations by bundling several transfers into a single transaction and reusing addresses and contracts.
+This behavior allows us to link attack transfers through the “guilt-by-association” heuristic [44].
 After removing transactions from unrelated bots that occasionally copy poisoning transactions from attackers, we can establish the existence of a few large distinct attack groups which we use to model the economics of address poisoning.
@@
 Despite high variability, most large groups manage to stay profitable by succeeding often, while smaller groups sometimes suffer from loss (by losing to the competition).
-There is no apparent early mover advantage.
 We also analyze attack strategies.
@@
 Based on our benchmarks, one sophisticated attack group appears to be using GPUs while others seem to rely on CPUs.
-Finally, we show that attackers are not as efficient as they could be, suggesting that the attack would be even more powerful if they optimized their strategies.
-This motivates us to propose a few mitigations for address poisoning attacks at the protocol, contract, wallet, and user levels.
+Finally, we propose a few mitigations for address poisoning attacks at the protocol, contract, wallet, and user levels.
+Compared to contemporary studies [47, 10], our paper offers the following contributions.
+First, our work detects attacks across multiple chains (BSC, Ethereum); in doing so, it identifies 13 times more attacks than prior work, and uncovers cross-chain attacks.
+Second, we devise improved clustering techniques, and use them to model group profitability and competition.
+Third, we perform novel experiments (e.g., victim profiling, success conditions, address generation modeling) to provide deeper insights into attack strategies and attacker computational capability.
 2 Background
@@
 Ethereum (and derived blockchains) features two types of addresses: wallet addresses (also called External Owned Accounts—EOA) and contract addresses.
-Wallet addresses to send Ether (ETH, the native currency) or interact with smart contracts (whose addresses are contract addresses).
+Wallet addresses are to send Ether (ETH, the native currency) or interact with smart contracts (whose addresses are contract addresses).
 Smart contracts are executed by all nodes using the Ethereum Virtual Machine (EVM).
-Smart contracts are developed in a higher level programming language (e.g., Solidity), and then compiled to “bytecode” intelligible by the EVM before being deployed.
+Smart contracts are developed in a higher-level programming language (e.g., Solidity), and then compiled to “bytecode” intelligible by the EVM before being deployed.
 Smart contracts allow users to perform actions, such as token implementation, token transfer.
 Of particular interest, ERC-20 tokens (e.g., USDT, USDC) are fungible assets, defined by the ERC-20 framework [41], for which creators specify a token name, symbol, and value unit.
-Essentially ERC-20 tokens allow developers to implement other cryptocurrencies or payment systems on top of Ethereum.
+Essentially, ERC-20 tokens allow developers to implement other cryptocurrencies or payment systems on top of Ethereum.
 For every action (e.g., sending ETH or ERC-20 tokens, or more generally, interacting with any contract), users pay “gas,” that is, transaction fees dependent on the computational complexity of the action requested.
@@
 (2) Apply elliptic curve (ECDSA) multiplication to produce a public key $K=k\times G$ where $G$ is the fixed generator point on the elliptic curve.
-(3) Hash (Keccak 256) the public key $K$, take the last 20 bytes, and add prefix $0x$ to obtain the address.
+(3) Hash (Keccak 256) the public key $K$, take the last 20 bytes, and add the prefix $0x$ to obtain the address.
 Elliptic curve (ECDSA) multiplication is computationally hard to invert, meaning that it is impractical to derive a private key from a public key.
-So, to generate a public key with a certain format (e.g., a vanity address), the only viable strategy is to perform a brute-force computation over potentially many private keys.
+To generate a public key with a certain format (e.g., a vanity address), the only viable strategy is to perform a brute-force computation over many private keys.
 The complexity of the brute force computation depends on the number of characters that have to be fixed in the desired public key address.
-For instance, it is roughly $16^{3}$ times harder to generate an (hexadecimal) address with a 3-character desired prefix compared to a random public key address.
+For instance, it is roughly $16^{3}$ times harder to generate an (hexadecimal) address with a 3-character desired prefix compared to a random address.
 Attackers adopt this brute-force technique to generate lookalike addresses.
@@
 Domain-name-based phishing attacks may resort to indirect monetization, e.g., ad impressions, or affiliate program enrollment [13, 21, 2].
-Blockchain address poisoning is markedly different from domain-name based phishing attacks in several respects.
-First, it focuses on the direct extraction of funds, and it may be more challenging for a victim to recover funds due to the irreversibility of cryptocurrency payments.
+Blockchain address poisoning is markedly different from domain-name-based phishing attacks in several respects.
+First, domain-name based attacks often proceed in multiple stages to steal funds: the victim first clicks on a malicious URL, is directed to a phishing website and is prompted to input sensitive information that can eventually be monetized [22].
+On the other hand, blockchain address poisoning only requires a single mistake from the victim—erroneously selecting an attacker address instead of a legitimate one.
+Furthermore, fund recovery is far more challenging due to the irreversibility of cryptocurrency transactions.
 Second, blockchain address poisoning attempts are permanently recorded in a public ledger, allowing for more accurate estimates of the losses [19].
@@
 We instead use the number of matching hexadecimal digits at the beginning and at the end.
-Fourth, domain-name based attacks often proceed in multiple stages to steal funds: the victim first clicks on a malicious URL, is directed to a phishing website and is prompted to input sensitive information that can eventually be monetized.
-On the other hand, blockchain address poisoning is far more direct.
-Victims only need to make the single mistake of submitting transactions by selecting the attacker’s address.
-Finally, blockchain address poisoning has differing costs to the attacker.
+Finally, address poisoning has differing costs to the attacker.
 They only need to deploy attack smart contracts (which can be even recycled from others for free), and one poisoning transfer costs only about a dollar in Ethereum and a cent in BSC.
 Generating lookalike addresses however has a variable cost that can be adjusted by the attacker, ranging anywhere from a laptop CPU to data centers of GPUs.
-Domain-based phishing attacks, on the other hand, require registering/hijacking domains, crafting phishing websites, and/or buying tool kits [3], all of which is expensive and time-intensive.
+Domain-based phishing attacks, on the other hand, require registering/hijacking domains, crafting phishing websites, and/or buying tool kits [3].
 3 The attack
@@
 We assume the existence of an attacker that passively monitors all public blockchains simultaneously in real-time.
-When available the attacker may also observe a state that has yet to be finalized in a chain such as transactions in the mempool.
+When available the attacker may also observe a state that has yet to be finalized in a chain, such as transactions in the mempool.
 The attacker can also actively engage with any chain by launching contracts and sending transactions using many accounts, bounded only by the chain and economic constraints.
 Finally, the attacker can generate lookalike addresses at a cost consistent with (publicly known) state-of-the-art hardware.
-We do not assume that the attacker has any out-of-band information about or agency towards a victim other than what can be obtained through public blockchains.
-Therefore attackers cannot influence the wallet software a victim is using, or make predictions about a victim’s transfers beyond what can be inferred through public signals.
+We do not assume that the attacker has any out-of-band information about or agency towards a victim, other than what can be obtained through public blockchains.
+Therefore attackers cannot influence the wallet software a victim is using, or make predictions about a victim’s transfers beyond what can be inferred through public signals (i.e., they cannot tailor the attack to each specific victim).
 We also assume that the security models for the individual chains hold, so attackers cannot, for example, perform denial of service on a chain or drop/modify a victim’s transactions.
@@
 In a zero-value transfer (Fig.
-1(b)), the attacker creates a transfer from $V$ to $L$ for the same token that $V$ sent to $R$, but with the value of zero.
+1(b)), the attacker creates a transfer from $V$ to $L$ for the same token that $V$ sent to $R$, but with a value of zero.
 Typically, token transfers can only be sender-initiated.
@@
 Wallets are represented by rectangles and contracts (or ERC-20 tokens created by contracts) are represented by ovals.
-We color attackers’s instances in pink.
+We color attackers’ instances in pink.
 First, an attack account ($A$) deploys a tailor-made attack contract ($AC$) and calls a function from it to execute multiple poisoning transfers at once.
 The attacker can also directly interact with the token contract to initiate a single poisoning transfer (without $AC$), which we observe in some cases despite its higher gas cost.
-The transaction $TX$ bundles together several potentially unrelated phishing transfers, in this case, the transfer of a lookalike token from $V1$ to $L1$ and a zero-value transfer from $V2$ to $L2$.
-Notice also that attacker $A1$ recycles their contract $AC1$ and $CT1$ to perform additional phishing transfers in $TX3$, and the infrastructure that $A1$ has embedded on a chain is typically distinct from that of other attackers such as $A2$’s use of $AC2$.
-It is also common for multiple attackers to target the same victim (e.g., $V1=V4$ ).
-As we will later describe, we have observed cases where attackers launch more than 100 poisoning transfers in a single transaction.
+The transaction $TX$ bundles together several potentially unrelated phishing transfers, in this case, the transfer of a counterfeit token from $V1$ to $L1$ and a zero-value transfer from $V2$ to $L2$.
+Notice also that attacker $A1$ recycles their contract $AC1$ and $CT1$ to perform $TX3$, and the infrastructure that $A1$ has embedded on a chain is typically distinct from that of other attackers, such as $A2$’s use of $AC2$.
+It is also common for multiple attackers to target the same victim (e.g., $V1=V4$ ), as we will later describe.
+We have observed cases where attackers launch more than 100 poisoning transfers in a single transaction.
 4 Detecting attacks
@@
 Guan and Li [10] capture all three poisoning transfers but follow an approach slightly different from ours.
-They flag suspicious transfers based on attack characteristics first, whereas we filter transfers when a victim interacts with two improbably similar, but different addresses.
-Our approach appears more robust to attacker manipulation (refer to Appendix A).
+While both filter transfers when a victim interacts with two improbably similar, our approach appears more robust to attacker manipulation (refer to §5.2 and Appendix B).
 We deploy our own Ethereum and BSC full nodes to scan the data efficiently instead of relying on third-party RPC endpoints.
@@
 We validate this choice of window size in §5.2.
-We look at all transfers in stablecoins – i.e., we search all transfers sent to (or from) addresses in $\mathcal{V}$ for potential tiny (or zero value) transfers.
+We look at all transfers [^2] We capture all transfers that emit a “transfer” signature, including any contract actions.
+in stablecoins – i.e., we search all transfers sent to (or from) addresses in $\mathcal{V}$ for potential tiny (or zero value) transfers.
 We also retrieve all ERC-20 token transfers for which an address in $\mathcal{V}$ is a source for potential counterfeit token transfers.
@@
 For each victim $V$, the corresponding intended address $R$, and the associated lookalike address $L$, we extract all related token transfer events collected in step 2.
-In other words, we only look at the cases where the victim $V$ interacts with two very similar addresses $R$ and $L$, making false positives statistically unlikely (except for those caused by typing mistakes, which we separately discuss in Appendix C).
+In other words, we only look at the cases where the victim $V$ interacts with two very similar addresses $R$ and $L$, making false positives statistically unlikely (except for those caused by typing mistakes, which we separately discuss in Appendix D).
 To identify counterfeit tokens, we acquire a list of ERC-20 and BEP-20 tokens from two sources: 1) Etherscan and BSCscan [^5] Tokens with at least a “Neutral” or “OK” reputation from Etherscan: https://etherscan.io/tokens and BSCscan: https://bscscan.com/tokens as of August 13, 2024.
@@
 We set the threshold $v_{0}$ to 10 USD to exclude unreasonably high transfers from the attacker.
-[^6] More than 98.79% of tiny transfers below 10 USD are less than 3 USD.
-We manually found some rare cases where the attackers sent nearly 10 USD as a tiny transfer.
+[^7] While 98.79% of tiny transfers ($<10$ USD) are less than 3 USD, we manually confirmed some rare tiny transfers were almost 10 USD.
     2. 2.
@@
 Recall that because the token was created by the attacker, they could implement any semantics they desire, notably having anybody send tokens to anybody else, without any authentication.
-  • A transfer is a payoff transfer if the victim $V$ sends an authentic token $AT$ ($v>0$) transfer to a lookalike address $L$.
-This is the transfer that ultimately monetizes the attack for the adversary.
+  • A transfer is a payoff transfer if a victim $V$ sends an authentic token $AT$ ($v>0$) to a lookalike address $L$, which ultimately monetizes the attack for the adversary.
 The top-right table in Figure 3 summarizes our labeling scheme.
-Red cells are attackers.
-Note that we convert the transferred value into USD on the day of the transfer.
-The price data is from CoinGecko API based on a contract address.
+Pink cells are attackers.
+Note that we convert the transferred value into USD on the transfer date.
+The price data is from the CoinGecko API based on a contract address.
 Step 5: Verifying payoffs.
@@
 First, the victim could have legitimately interacted with two similar addresses by sheer coincidence.
-This is statistically unlikely, specifically because we already filter out addresses particularly vulnerable to the “birthday paradox” (refer to Appendix B for our detailed analysis).
+This is statistically unlikely, specifically because we already filter out addresses particularly vulnerable to the “birthday paradox” (refer to Appendix C for our detailed analysis).
 Second, the attack may have taken place over a time interval exceeding our 20-minute monitoring window.
@@
 If 1) the destination address never sent any tokens, i.e., the attacker does not redeem the payoff transfer, and 2) $d$ is more than 20, i.e., computationally infeasible to generate such a $L$, we assume that the destination address was actually a typo.
-We indeed find 681 accidental transfers from 501 users, totaling 5.5 million USD in losses (almost all in Ethereum, with only 57,000 USD in BSC): Appendix C for more detailed method, results, and analysis.
+We indeed find 681 accidental transfers from 501 users, totaling 5.5 million USD in losses (almost all in Ethereum, with only 57,000 USD in BSC): Appendix D for more detailed method, results, and analysis.
 4.2 Clustering attackers
@@
 Each address poisoning attack consists of 1) a poisoning transfer ($TR$), 2) a poisoning transaction ($TX$) that initiates $TR$, 3) a lookalike wallet address ($L$), 4) a counterfeit token ($CT$), 5) an attack contract ($AC$) that initiates $TX$, and 6) an attack wallet address ($A$) that interacts with $AC$ (see Figure 2).
-We associate each poisoning transfer $TR$ with the corresponding attack instance $\{TX,L,CT,AC,A\}$.
-Similar to the “guilt-by-association” technique [43], we merge two attack sets if they share at least one of $TX$, $L$, or $A$.
+We associate each poisoning transfer $TR$ with the corresponding attack “transfer set” $\{TX,L,CT,AC,A\}$.
+Similar to the “guilt-by-association” technique [44], we merge two attack transfer sets if they share at least one of $TX$, $L$, or $A$.
 In other words, we assume that 1) two poisoning transfers in the same transaction belong to the same attacker, 2) an attacker is economically rational and specifies addresses $L$ they control (to successfully receive payoffs), and 3) an attacker does not share the private key of the attacker’s wallet $A$ with other attackers.
@@
 For instance, in Figure 2, we 1) merge ($TR_{1}$, $TR_{2}$) because both transfers are in the same transaction, 2) merge ($TR_{1}$, $TR_{3}$) if $L_{1}=L_{3}$, and 3) merge ($TR_{3}$, $TR_{4}$, $TR_{5}$, $TR_{6}$) if $A_{1}=A_{2}$.
+We keep merging attack transfer sets until there is no overlap of $TX$, $L$, and $A$ between attack groups.
 Having applied this process to our entire dataset, we notice that two seemingly distinct large attack groups—active at different times or using different attack strategies—can be merged due to a few overlapping attack addresses $A$.
 Manual investigation revealed that those accounts appear to be bots copying poisoning transactions from two different attack groups with little or no modification, leading to a seemingly erroneous merge.
-Indeed, Guan and Li [10] performed similar clustering techniques, but did not consider the existence of bots and appeared to form one large cluster.
+Guan and Li [10] perform similar clustering techniques, but did not consider the existence of bots and appeared to form one large cluster.
 To prevent such issues, we explored several methods to exclude transactions from bots.
@@
 We then set a threshold to exclude addresses that have a low attack ratio.
-We document the details of those bots’ behavior and the bot exclusion process in Appendix D.
+We document the details of those bots’ behavior and the bot exclusion process in Appendix E.
 5 Results and evaluation
@@
 5.1 Detection results
-We present the summary statistics for poisoning transfers in Table 1 and confirmed payoff transfers in Table 2
+We present the summary statistics for poisoning transfers in Table 1 and confirmed payoff transfers in Table 2.
 Table 1: Attack summary statistics
@@
 Max Std.
-dev Count Victims (USD) (USD) (USD) (USD) (USD) (USD) Ethereum 79,344,412 0.001 2,160 45,653 20,000,000 521,663 1,738 1,502 BSC 4,490,804 0.000 85 1,164 279,489 8,366 4,895 4,004
+dev Count Victims (USD) (USD) (USD) (USD) (USD) (USD) Ethereum 79,344,412 0.001 2,160 45,653 20,000,000 514,142 1,738 1,502 BSC 4,490,804 0.000 85 1,164 279,489 7681 4,895 4,004
 On Ethereum, we find 17.3M poisoning transfers: 300,000 tiny transfers, 7.2M zero-value transfers, and 9.9M counterfeit-token transfers over 1.7M transactions.
@@
 Attackers on Ethereum receive 1,738 transfers from 1,502 victims, totaling nearly 80M USD (avg.
-45,853 USD, median 2,169 USD, with high variance), with a success rate is 0.01%8880.13% spam click-through rate in Twitter [9].
+45,853 USD, median 2,169 USD, with high variance), with a success rate is 0.01%9990.13% spam click-through rate in Twitter [9].
 Some victims sent assets to the attacker’s address(es) multiple times.
@@
 5.2 Detection evaluation
-Manual evaluation on successful cases.
+Evaluation with external dataset.
+For our detector evaluation, we obtained from Guan and Li [10] their evaluation dataset and detection results.
+[^10] The authors kindly made the evaluation open source recently: https://github.com/DS2L/Poison-Hunter.
+The data consists of 5,890 lookalike addresses and 1,154 benign (non-lookalike) addresses.
+The authors include lookalike addresses only when two sources (Forta and Etherscan) report them and label DeFi entities as benign based on the Etherscan Label Cloud.
+[^11] https://etherscan.io/labelcloud We only use our results from Nov.
+2022 to Feb.
+2024 to align with their experiment.
+We discovered and manually confirmed that 45 addresses are mislabeled as lookalike—there are no corresponding poisoning transfers.
+We have contacted Guan and Li, who agreed to remove these mislabeled entries from their dataset.
+After removing mislabeled entries, our detector achieves 97.7% accuracy, 100% precision, 97.2% recall, and a 98.58% F1-score.
+In absolute terms, we identify 5,681 lookalike addresses with 164 false negatives and 0 false positives.
+We investigate those 164 false negatives using Etherscan.
+A full discussion of the whole analysis can be found in Appendix A.
+Here we summarize the high-level results.
+In 87 instances of false negatives, we fail to identify the lookalike address because the victim engaged in a benign transfer that resembles a poisoning transfer.
+For instance, the victim may have sent a legitimate zero-value transfer to the intended address for testing.
+Our detector cannot distinguish between that case and that of an attacker creating a zero-value transfer appearing to originate from the victim, and sent to a lookalike address.
+In such a case, our detector conservatively does not classify the event as an attack, and tags the corresponding lookalike address as legitimate.
+23 false negatives are due to attacks being outside our time-window range (i.e., attackers target the victim after 20 minutes).
+16 false negatives come from attackers targeting other tokens (e.g., TrueUSD, BUSD) while we only look at major stablecoin transfers.
+The rest (38 false negatives) can be primarily attributed to data collection quirks.
+Importantly, our choice of similarity threshold $(3,4)$ or that our detector only considering senders as potential victims is not the cause of any of the false negatives.
+Guan and Li’s detector discovers 144 new addresses we did not catch, while our detector captures 96 addresses that Guan and Li’s did not.
+Guan and Li’s detector employs additional checks using attack characteristics (e.g., checking a transaction initiator) and filters out benign transfers that resemble poisoning transfers (i.e., mitigating the first case of false negatives we identified above).
+However, Guan and Li’s approach (slightly overfitting to known attacks) may miss poisoning transfers when the attacker modifies their attack strategy (e.g., using a lookalike address to launch a poisoning transaction).
+Manual evaluation on successful cases and attack contracts.
 Statistically, the probability of getting a false positive is exceedingly low because 1) we only look at transfers where the victim has interacted with two similar addresses, and 2) we observe poisoning transfers between an intended and a payoff transfer take place.
@@
 We sorted payoff transfers by decreasing the amount and checked the 30 largest transfers for both chains.
-We visit each victim’s page on external blockchain scanning websites (e.g., Etherscan and BSCscan) and manually confirmed 1) the intended (original) transfer, 2) the poisoning transfer, and 3) the payoff transfer.
+We visited each victim’s page on external blockchain scanning websites, Etherscan and BSCscan, and manually confirmed 1) the intended (original) transfer, 2) the poisoning transfer, and 3) the payoff transfer and their order.
 We verified that all 30 cases are indeed payoff transfers for both Ethereum and BSC (i.e., no false positives).
 Those 30 cases account for 53.5M (67%) of Ethereum losses and 2M (45%) of BSC losses.
-For accountability purposes, we list the top 30 successful cases for Ethereum and BSC in Appendix G.
-Comparison with previous work.
-We directly compare our detection results by focusing on 12 and 16 months of data to align with Ye et al.’s work [46] and Guan and Li’s [10] work, respectively.
-We detect more payoff transfers than Ye et al.
-We find 1,059 successful cases with 36.6M USD in losses while Ye et al.
-found 560 successful cases with at most 28M USD in losses, including two separate (non-address poisoning) attacks they investigate.
-In terms of poisoning transfers, we identify 4.9M zero-value token transfers, compared to 21M for Ye et al.
-It is unclear whether the difference comes from 1) false positives on Ye et al.’s side, (i.e., benign zero-value token transfers), or 2) false negatives on our side.
-We find a similar total loss on Ethereum (during the same period) as Guan and Li [10], but we detect more than 2.2M counterfeit token transfers, mostly from tokens with modified symbols (refer to a more detailed comparison in Appendix A).
+We perform the same operation for the contracts used for poisoning transfers ($AC$).
+We sort $AC$ based on the number of poisoning transfers, manually look at the top 30 $AC$ for each chain on external blockchain scanning websites, and confirm at least one poisoning transfer (and its corresponding intended transfer on the victim’s page).
+We verify that all the contracts are indeed attack contracts, and they account for 4.9M (28%) on Ethereum and 210.1M (83%) on BSC.
 Checking the contract verification.
@@
 We exclude verified attack contracts and counterfeit token contracts from our clustering to avoid potential erroneous merge.
-Window range.
-To make the chain scanning process efficient, we assume that the attack happens in the next 20 minutes (100 blocks in Ethereum) after the original transfers.
-We evaluate this assumption by collecting data over 200,000 blocks starting from 16,950,603 (Apr.
-1st, 2023), using both a 100-block range and a 200-block range, and comparing results.
-Both ranges identify the same number of successful cases (90).
-The 200-block range only identifies two additional tiny transfers besides the 250 tiny transfers identified with a 100-block range.
-With a 100-block range, we find 102,154 zero-value transfers; extending to 200 blocks only identifies an additional 27 zero transfers.
-Finally, a 100-block window captures 704,998 counterfeit token transfers, whereas a 200-block window brings this up to 706,332.
-In short, doubling the window size (20 minutes to 40 minutes) does not dramatically increase the number of attacks we discover.
+Parameter selection.
+We set key assumptions about the window size (100 blocks) and the similarity threshold (3, 4) to make the chain scanning process efficient and reduce false positives.
+We evaluate those assumptions by collecting data over 200,000 blocks starting from 16,950,603 (Apr.
+1st, 2023), and performing a detection with a window size of 200 and similarity score of (3, 3).
+Table 3 compares the different parameters and summarizes the results on the number of poisoning transfers, lookalike addresses, and payoff transfers.
+The 200-block range captures 1,363 (0.17%) more poisoning transfers, 1,139 (0.27%) more lookalike addresses, and the same number of payoff transfers.
+Hence, doubling the window size does not dramatically increase the number of attacks we discover.
 We believe that collecting other transfers within the same poisoning transaction (§4.1: Step 2) has significantly helped the coverage.
+When we select the threshold of (3, 3) instead of (3, 4), we capture 2,937 (0.36%) additional poisoning transfers and 972 more lookalike addresses (0.23%).
+However, if we decrease the threshold by 1, we are $\approx\sqrt{16}=4$ times more likely to find benign lookalike addresses.
+One additional payoff transfer identified with a threshold of (3, 3) was indeed a false positive.
+In short, changing those parameters does not appear to outweigh the cost of data collection and false positives.
+Table 3: Detection results with different parameters. T: tiny, Z: zero-value, C: counterfeit token transfer, L: lookalike address
+Window size Similarity Poisoning (T, Z, C) L Payoff
+100 (3, 4) 807,402 (250, 102,154, 704,998) 427,317 90
+200 (3, 3) 808,765 (252, 102,181, 706,332) 428,456 90
+100 (3, 3) 810,339 (265, 102,363, 707,711) 428,289 91
 5.3 Clustering results
@@
 We choose a threshold of 0.5, which 1) avoids erroneous merge of large clusters and 2) only removes 1.49% of all transfers.
-We provide a detailed explanation of this threshold selection and demonstrate the robustness of our clustering results (through temporal clustering) in Appendix D.
-Table 3 describes the number of attack instances and assigns Group IDs sorted by the number of lookalike addresses in each group on Ethereum.
-The notation follows §3.
-We only list the top eight biggest groups because the size of the eighth group is significantly smaller than the top seven (i.e., less than 100,000 lookalike addresses with a few successful cases).
-Different groups exhibit different attack strategies; not all groups use three poisoning transfers.
+We provide a detailed explanation of this threshold selection in Appendix E.
+In total, we identify 49 groups (with more than one lookalike address), corresponding to 97.4% of poisoning transactions, corroborating our detection results.
+Table 4 describes the number of attack instances and assigns Group IDs sorted by the number of lookalike addresses in each group on Ethereum, following the notation in §3.
+We only list the top eight groups because the eighth group is significantly smaller than the top seven (i.e., less than 100,000 lookalike addresses with a few successful cases).
+Different groups exhibit different attack strategies; not all groups use all three poisoning types.
 Only Group 1 uses tiny transfers.
-Groups 3, 4, and 7 do not have any counterfeit token contracts, $CT$, and thus do not use any counterfeit-token transfers, and instead focus on zero-value transfers.
-We additionally collect contract bytecode from our Ethereum node and list the number of distinct contracts in a parenthesis.
-In Group 2, for example, all $CT$ and $AC$ are identical, indicating the reuse of the same source code (but with different addresses).
+Groups 3, 4, and 7 do not use any counterfeit token transfers, and instead focus on zero-value transfers.
+Additionally, we collect contract bytecode from our Ethereum node and list the number of distinct contracts in parentheses.
+In Group 2, for example, all $CT$ and $AC$ are identical, indicating the same source code reuse (but with different addresses).
 Given that the contract bytecode changes even with a slight alteration in the source code, the high number of identical contracts corroborates the validity of our clustering techniques (especially since our clustering does not rely on those contracts).
-Table 3: Attack group general statistics. Notations follow Fig. 2). The distinct number of contracts is in parentheses.
+Table 4: Attack group statistics. Notations follow Fig. 2. The distinct number of contracts is in parentheses.
 Group $\mathbf{L}$ $\mathbf{CT}$ $\mathbf{A}$ $\mathbf{AC}$ $\mathbf{R}$ $\mathbf{TR}$ $\mathbf{TX}$ 1 1,390,902 3,309 (517) 4,818 2,147 (67) 1,091,130 6,485,270 640,600 2 1,221,956 1,491 (1) 382 259 (1) 970,832 3,329,931 573,500 3 1,199,268 0 (0) 241 242 (2) 833,140 2,317,361 127,215 4 473,140 0 (0) 27 27 (4) 308,079 595,720 38,675 5 465,389 392 (149) 241 227 (74) 375,444 1,003,463 15,222 6 423,107 81 (46) 104 13 (13) 280,045 1,303,112 63,856 7 272,557 0 (0) 27 39 (39) 272,559 600,876 16,368 8 99,020 43 (11) 399 3 (2) 75,742 210,831 13,589
@@
 2023.
+While several top groups stop operations around May 2024 (in Figure 4, Ethereum), the attack is still active today; we observe a drastic increase in BSC in June 2024 (in §5.1).
 As we will see in Table 5, we could not confirm an early mover advantage (i.e., later comers are still profitable).
 The variation in the attack strategies and the active periods across attack groups further supports our clustering results.
+All in all, we consistently see behavioral heterogeneity across the groups we identified, such as distinct attack strategies, identical contract reuse, and different activity periods.
+In Appendix E, we also demonstrate the robustness of our clustering results through temporal clustering; most attack groups stay stable/distinct over time.
 Figure 4: Number of poisoning transfers from each attack group over time (weekly basis)
@@
 6.1 Targeted population
-We first look at victim profiles, so that we can better understand who the attackers target, and eventually propose defenses.
+Attackers appear to select their targets in advance because the attack typically happens right after the target’s original transfer (mostly within 20 minutes, as shown in §5.2).
+We further investigate whether attackers disproportionately target certain populations compared to the average stablecoin user.
 To make the attack more profitable, we conjecture that attackers are more likely to target rich addresses.
@@
 For each victim, we retrieve the sum of their three stablecoins (USDT, USDC, DAI) balance just before when the victim receives the first poisoning transfer from each attack group.
-We also create a baseline group by randomly sampling 1,000 USDT users and looking up their balance before the most recent transfer.
+We also create a baseline group by randomly sampling 1,000 USDT users and looking up their balance (for three stablecoins) before the most recent transfer.
 Figure 5 illustrates the cumulative distribution function (CDF) of the total stablecoin balance for the victims in each of the seven attack groups and the baseline USDT users.
@@
 First, we discover that the attacker tends to imitate addresses ($\mathcal{R}$) from centralized entities.
-By manually looking at the list of addresses in $\mathcal{R}$ on Ethereum that the attacker imitates to generate $L$ the most, we find that those addresses often belong to centralized exchanges (e.g., Coinbase, Binance, Bybit) probably because they are one of the most active accounts.
+By manually looking at the list of addresses in $\mathcal{R}$ on Ethereum that the attacker imitates the most, we find that those addresses often belong to centralized exchanges (e.g., Coinbase, Binance, Bybit), probably because they are one of the most active accounts.
 However, most addresses are hot wallets, with which normal users typically do not interact.
@@
 This appears to be a waste of time and money, as excluding such addresses would reduce address generation and transaction costs.
-We have more detailed information in Appendix E.
+We have more detailed information in Appendix F.
 Second, we observe distinctive patterns in generating lookalike addresses.
@@
 (b) Other groups
-Figure 7: Number of lookalike addresses for which the first $x$- and the last $y$-digit match the intended recipient address.
+Figure 7: Number of lookalike addresses for which the first $x$- and the last $y$-digit match the intended address.
 Figure 8 depicts the distributions of lookalike address similarity for the top four groups.
@@
 The difference in distributions also suggests that different attack groups may have employed different infrastructures, which we will investigate in §7.2.
-Figure 8: Distributions of the lookalike address similarity (for the top 4 attack groups): y-axis is on a log scale
+Figure 8: Distributions of the lookalike address similarity (for the top 4 attack groups): y-axis is on a log scale.
 Finally, we investigate whether any attack group re-uses lookalike addresses across different chains.
@@
 We next look into which conditions are satisfied for an attack to be successful.
-We test our hypothesis that the more similar the lookalike address is to the intended recipient address, the more likely the victim will make a mistake.
+We test our hypothesis that the more similar the lookalike address is to the intended address, the more likely the victim will make a mistake.
 In Figure 9, the bubble size represents the success ratio (the number of successes to the number of lookalike addresses) depending on similarity.
@@
 This result suggests that the attack may have been more profitable to generate a larger number of slightly less convincing lookalike addresses and conduct more attacks.
-Figure 9: Success probability and address similarity. The first $x$- and the last $y$-digit match the intended recipient address.
-We next focus on successful cases and examine how address similarity (or attack timing) relates to success.
-Figure 10 (left) is the number of lookalike addresses targeting each victim.
+Figure 9: Success probability and address similarity. The first $x$- and the last $y$-digit match the intended address.
+We next focus on successful cases and examine how address similarity or attack timing relates to success.
+Figure 10 (left) shows the number of lookalike addresses targeting each victim.
 In 79% of the successful attacks on Ethereum, multiple lookalike addresses target one victim, indicating competition among attackers.
-This competition allows us to investigate how a victim selects one particular lookalike address (i.e., a winner $L_{\text{win}}$) over a pool of other lookalike addresses (i.e., victims’ perception).
+This competition allows us to investigate how each victim selects one particular lookalike address, a winner $L_{\text{win}}$, over a pool of other lookalike addresses, while partially controlling for individual user attributes.
 We focus on two aspects: how similar the winner $L_{\text{win}}$ is to $R$; and how early the winning address appears, compared to other lookalike addresses.
 For each successful attack, we extract the list of corresponding lookalike addresses $\mathcal{L}$, then check how $L_{\text{win}}$—the recipient of the payoff transfer—ranks in these two metrics.
-A smaller rank implies a higher similarity; and an earlier attack, respectively.
+A smaller rank implies a higher similarity and an earlier attack, respectively.
 Figure 10 shows that, while the number of competitors (left) decreases somewhat linearly, a disproportionately large proportion of the most similar addresses (center), and/or early-comers (right) tend to win the competition.
@@
 While those results provide insights from chain activities, the success also relies on what wallets or chain scanning services victims use.
-Appendix 8 compares how poisoning attacks appear in those services’ UI and qualitatively discusses when and where victims are most likely to fall for the attack.
+Appendix G compares how poisoning attacks appear in those services’ UI and qualitatively discusses when and where victims are most likely to fall for the attack.
 Figure 10: Histogram for the number of lookalike addresses per victim (left), the ranking of the $L_{\text{win}}$ in terms of similarity (middle), and timing (right).
@@
 Despite the high variance in payoff amounts as shown in Table 2, those groups manage to amortize the revenue across many payoff transfers while some small groups do not.
-Given the attack scale (i.e., Group 1 has invested over 8 million USD), some groups appear to be well-funded cyber-crime entities.
+Given the attack scale (i.e., Group 1 has invested over 8 million USD), some groups appear to be well-funded cybercrime entities.
 However, none of the attack instances comes from any sanctioned entities in the SDN list [42].
@@
 7.1 Mathematical formulation
-We assume that the attacker wants to match the first $a$ and the last $b$ digits of the intended recipient and maximize the total number of digits $a+b=d$, with no preference for $a$ over $b$.
+We assume that the attacker wants to match the first $a$ and the last $b$ digits of the intended address and maximize $a+b=d$, with no preference for $a$ over $b$.
 For instance, the attacker is indifferent between $(a,b)=(3,6)$ or $(4,5)$ as long as $(a\geq 3,b\geq 4)$ holds, per §4.1.
-The attacker has the set of target addresses ($\mathcal{R}$) they want to imitate.
+The attacker has a set of target addresses ($\mathcal{R}$) they want to imitate as discussed in §6.1.
 We assume that the probability of generating a given address follows a uniform random distribution over the entire address space, and that the number of target addresses $|\mathcal{R}|$ is much smaller than the total address space ($16^{40}$).
 If $|\mathcal{R}|=r$, the probability of generating a target address is
-$p=1-\left(\frac{16^{d}-1}{16^{d}}\right)^{r}\ ,$ (1)
-which reduces to $p=1/16^{d}=2^{-4d}$ when $r=1$ (i.e., a single address is targeted).
-The probability of finding a collision after $k$ trials follows the geometric distribution $P(X=k)=p\cdot(1-p)^{k-1}$, and the expected number of trials for a collision is $E[X]=1/p$.
-Using a first-order Taylor approximation in Eqn.
-(1) yields $p\approx r/16^{d}$, and the resulting expected number of trials is $E[X]\approx 2^{4d}/r$.
-In short, the collision probability linearly increases with the size of the target address space $r$, and, accordingly, the expected number of trials linearly decreases with $r$.
+$p=1-\left(\frac{16^{d}-1}{16^{d}}\right)^{r}\ .$ (1)
+Using a first-order Taylor approximation yields $p\approx r/16^{d}$.
+The probability of finding a collision after $k$ trials follows the geometric distribution $P(X=k)=p\cdot(1-p)^{k-1}$, and the expected number of trials for a collision is $E[X]=1/p\approx 2^{4d}/r$.
+In short, the expected number of trials exponentially increases with the the number of matched digits, and linearly decreases with the size of the target address space.
 7.2 Simulating address generation
@@
 This allows us to measure the number of addresses per second (APS) that can be generated, which in turn gives us a lower bound on the computational resources attackers must possess to produce the lookalike addresses we observed in the wild.
-We implement 1) a naive version using Python with the web3py [^10] https://github.com/ethereum/web3.py library; and 2) an optimized version leveraging CPU multiprocessing using Python with the multiprocessing, coincurve, [^11] https://github.com/ofek/coincurve, Python bindings for a heavily optimized C library for elliptic curve operations.
-and pycryptodome [^12] https://github.com/Legrandin/pycryptodome/, the default hashing package used by web3py.
-libraries.
-In addition, we use two popular vanity address-generating software: Vanity-ETH, [^13] https://vanity-eth.tk/ a Javascript-based web application, and Profanity2, [^14] https://github.com/1inch/profanity2 a C++-based GPU implementation.
+For address generation software, we have tested several combinations of libraries: web3py [^13] https://github.com/ethereum/web3.py., coincurve [^14] https://ofek.dev/coincurve/, a set of Python bindings for a heavily optimized C library for elliptic curve operations., pycryptodome [^15] https://github.com/Legrandin/pycryptodome/, the default hashing package used by web3py., with/without concurrency, and multiprocessing.
+We also implement the address generation scripts in Go (with concurrency), which yields results similar to those obtained with Python.
+We additionally use two popular vanity address software: Vanity-ETH, [^16] https://vanity-eth.tk/ a Javascript-based web application, and Profanity2, [^17] https://github.com/1inch/profanity2 a C++-based GPU implementation.
 We elect, for simplicity, not to check whether the generated address is in the set of target addresses $\mathcal{R}$, which will produce an optimistic bound on performance, and thus a conservative bound on the attackers’ implied hardware capabilities.
-We test our implementations on two hardware configurations, a Macbook Air (OS Version 13.1) with an 8-core Apple M1 base chip; and a server workstation (Ubuntu 22.04 LTS) with an Intel(R) Xeon(R) Silver 4214 CPU and an Nvidia RTX 8000 GPU.
+We test our implementations on two hardware configurations, a MacBook Air (OS Version 13.1) with an 8-core Apple M1 base chip, and a server workstation (Ubuntu 22.04 LTS) with an Intel(R) Xeon(R) Silver 4214 CPU and an Nvidia RTX 8000 GPU.
+Other hardware settings (e.g., different generations of Mac chips) only result in minor variations in outcomes.
 We run our simulations on the CPU on both machines and on the GPU on the second machine.
 We summarize our simulation results in Table 6.
-The naive Python implementation which models what might be an expected “first-attempt” by an attacker produces in the order of $10^{3}$$10^{4}$ addresses per second.
-The more optimized implementation produces an order of magnitude improvement ($10^{4}$$10^{5}$).
+We focus on 1) a naive implementation which models what might be an expected “first-attempt” by an attacker, using the Python web3py library, and 2) the best performing implementation: coincurve + pycryptodome + multiprocessing.
+The naive software implementation produces in the order of $10^{3}$$10^{4}$ addresses per second.
+The more optimized software implementation produces an order of magnitude improvement ($10^{4}$$10^{5}$).
 The Vanity-ETH implementation is much slower than the standalone clients.
 The CUDA implementation from Profanity2 is three orders of magnitude faster than our best CPU implementation with roughly half a billion APS.
-The Profanity2 developers self-report about a billion APS on similarly priced consumer hardware.
-However, the overhead imposed by checking set inclusion in $\mathcal{R}$ should be much higher in the GPU implementation.
+The Profanity2 developers self-report about a billion APS on similarly priced consumer hardware [^18] https://github.com/1inch/profanity2?tab=readme-ov-file#benchmarks---current-version.
 Table 6: Generated addresses per second (APS)
@@
 Based on the distribution of the matching digits $d$ for each attack group, we next attempt to infer the attacker’s hardware capability.
-We define two units: CPU-day and GPU-day, the number of addresses the attacker can generate in one day per machine with a CPU and a GPU, respectively.
+We define two units: CPU-day and GPU-day, the number of addresses attackers can generate in one day per machine with a CPU and a GPU, respectively.
 From the above benchmarks, 1 CPU-day is $3.98\times 10^{10}$ addresses and and 1 GPU-day is $4.46\times 10^{13}$, respectively.
@@
 We thus suspect they perform a substantial amount of their computation via GPUs.
-Group 1 could also have developed specialized hardware for address generation (e.g., ASICs), which would further reduce generation time.
 Other groups ($d=14$) only used 1.81 CPU-days, or $1.6\times 10^{-3}$ GPU-days.
@@
 For CPU, an hourly rate (on-demand) for r6i.4xlarge is \$1.008; \$24.19 for one CPU-day.
-Using these numbers, we estimate that Group 1 would have paid (at most) 1.7M USD to generate a 20-digit match, which is well below its 4M USD profit described in Table 4.
-Furthermore, these numbers are likely higher than the costs incurred by self-hosting hardware.
-For other groups working with $d=14$, CPU costs ($1.81\times 24.19\approx 43$ USD) are dominated by transaction costs.
+Using these numbers, we estimate that Group 1 would have paid (at most) 1.7M USD to generate a 20-digit match, which is below its 4M USD profit described in Table 5.
+Furthermore, these numbers are likely higher than the costs incurred by self-hosting machines.
+Notably, we do not consider specialized hardware for address generation (e.g., ASICs), which would further reduce the generation time/cost for Group 1.
+For other groups with $d=14$, CPU costs ($1.81\times 24.19\approx 43$ USD) are dominated by transaction costs.
 8 Mitigations
@@
 Protocol-level mitigations: One mitigation is to map human-readable strings to each address, like DNS for IP addresses.
-Some entities already have adopted domain names on wallet addresses through the Ethereum Domain Name System (ENS).
-Human-readable addresses would reduce address poisoning as well as accidental transfers (Appendix C) and enhance usability.
-ENS, however, fosters other security risks such as typosquatting, name hijacking (i.e., re-using the expired domain names) [44], or name collisions [14].
+Some entities already have adopted domain names for wallet addresses through Blockchain Domain Name System (BNS).
+Human-readable addresses would reduce address poisoning as well as accidental transfers (Appendix D) and enhance usability.
+BNS, however, fosters other security risks such as typosquatting [22], name hijacking (i.e., re-using the expired domain names) [45], or name collisions [14].
 Another protocol-level solution is to raise the cost of generating an address and make the attack economically less effective [30].
@@
 These additional costs could help render the attack model economically impractical.
-Contract-level mitigations: One solution is to modify the major stablecoin contracts and ask permission from the token sender even when the amount is zero.
-Function modification requires each ERC-20 token issuer to upgrade its smart contract despite limited incentives for developers.
-This solution only mediates zero-value transfers but not tiny and counterfeit token transfers.
-In addition, token owners can quickly blacklist the attacker’s address.
+Contract-level mitigations: One solution is to disallow zero-value transfers in “transferFrom,” despite the existence of use cases (e.g., testing by developers).
+Specifically, we could theoretically modify the major stablecoin contracts (if they are upgradable) and mandate permission from the token sender, even when the transfer amount is zero.
+However, the community [^20] https://github.com/OpenZeppelin/openzeppelin-contracts/issues/3931 has concluded that no changes to these contracts should be made, as removing zero-value transfers (in transferFrom) would break existing code (e.g., “_becomeImplementation” in CDaiDelegate.sol (Compound) that relies on such zero-value transfers.)
+Another potential (post-attack) solution is to quickly blacklist attack addresses.
 For the largest payoff transfer we observe, the Tether team (USDT) blocked the attacker’s address within an hour, which prevented the attacker from moving the retrieved funds to another address.
@@
 Xia et al.
-[43] identified around 10,000 counterfeit tokens on Uniswap (50% of all Uniswap tokens) through ground truth labels provided by the platform, address clustering, and machine learning-based techniques.
-They estimated these counterfeit tokens collectively led to financial losses of 16M USD from approximately 40,000 victims.
+[44] identified around 10,000 counterfeit tokens on Uniswap (50% of all Uniswap tokens) through ground truth labels provided by the platform, address clustering, and machine learning-based techniques, leading to a 16M USD loss from approximately 40,000 victims.
 He et al.
@@
 The authors identified 26,333 phishing websites with 3,486 phishing blockchain accounts.
+Muzammil et al.
+[22] studied typosquatting on BNS, and identified over 26,000 typosquatting names and thousands of transactions sent to the addresses mapped to these squatting names.
 Other studies on free giveaway scams reported 24M–69.9M USD losses [19], including 872K USD losses from Twitter [17].
-Based on these other works, blockchain address poisoning appears to be one of the largest (in terms of aggregate losses) and most wildly targeted criminal activity taking place on-chain.
-Closest to our efforts, Ye et al.
-[46] examined three types of blockchain scams, one of which is address poisoning on Ethereum with zero-value transfers.
-In parallel, and contemporary to our work, Guan and Li [10] also studied address poisoning on Ethereum and discuss all three poisoning transfers.
-We provide a detailed comparison with our method and results in §4.1, §5.2 and Appendix A.
-The Chainalysis’s blog post [34] conducted the case study (WBTC 68M loss) and analyzed its related addresses (victims, lookalike addresses).
+Based on these other works, blockchain address poisoning appears to be one of the largest (in terms of aggregate losses) and most widely targeted on-chain criminal activities.
+Finally, we clarify our contributions in contrast to other contemporary works, most notably Guan and Li [10] and Ye et al.
+[47].
+There are three major differences between our approach and Guan and Li’s.
+First, our clustering is more fine-grained.
+While we use guilt-by-association techniques to cluster attack instances (similar to theirs), crucially, we detect copying bots, which will erroneously merge distinct groups into one large cluster, as described in §4.2 and Appendix E.
+Second, we ran novel experiments and analyses, including victim characteristics, success conditions, group competitions, and address generation modeling, which provide deeper insights into attacker strategies and computational capability.
+Third, we uncovered 13 times more attacks (270M total) than prior works and compared the results between the two chains.
+Despite the similar chain structure, BSC requires more effort and optimization due to the large number of transfers.
+(Our BSC database is 8.6 times larger than Ethereum’s).
+That dataset also allows us to uncover cross-chain attacks.
+Another contemporary piece, by Ye et al.
+[47], examines zero-value transfers, but not other types of poisoning transfers, which we do cover.
+In Appendix 7, we adjust our measurement period to provide a direct comparison of detection outcomes between our work and those two studies.
+While most industry reports are case studies, a Chainalysis’s blog post [35] extended one case study (WBTC \$68M loss) to identify an additional 82,031 lookalike addresses.
+While commendable, this remains small compared to the whole ecosystem, as shown by our work.
 10 Conclusion
@@
 11 Ethical considerations
-The concept of blockchain address poisoning is already the subject of some online discussions, and, as this paper shows, this issue is currently extensively exploited in the field.
-As such, wallet operators and (some) users are most likely already aware of at least the theoretical possibility of the attack.
-What our work does, though, is to demonstrate its practical severity in economic terms – and we hope it will be a call to arms to act sooner rather than later.
-While attackers could potentially misuse some of our findings to improve their strategies, besides raising community awareness, the main contribution of our paper is to extensively describe the attack mechanisms.
+While the concept of address poisoning is already the subject of many online discussions, we notified Etherscan and the MetaMask security team of our findings; MetaMask acknowledged that they are aware of address poisoning.
+Although attackers could potentially misuse some of our findings to improve their strategies, the main contribution of our paper is to demonstrate its practical severity in economic terms.
+We hope that our work can help users recognize when they are in targeted populations, and accordingly, take precautionary measures.
 We also introduce possible protection strategies.
 Our proposed mitigations can hopefully inspire wallet operators.
-Likewise, we hope that our work can help users recognize when they are in targeted populations, and accordingly, take precautionary measures.
-Finally, Appendix G, specifically Table 9, includes full transaction hashes, which could potentially be used to re-identify the wallet addresses of the potential victims (whom we do not directly list in the table).
-We decided to present the table as is, as obfuscating part of the transaction hash would impede reproducibility or external validation of our findings, for little to no privacy benefit: users of Ethereum (or BSC) are relying on a public blockchain, in which all records are publicly available.
-Furthermore, wallet addresses are not, on their own, considered personally identificable information, as 1) there is no one-to-one mapping to specific individuals, as a given person may hold several wallets, and a single wallet may be shared by multiple people (e.g., belonging to the same organization), and 2) a wallet address does not, in itself, contain information directly mapping to a real-world identity.
+12 Open science
+We released the detection result dataset, including over 17 million attack attempts on Ethereum and successful payoff transfers.
+[^22] https://kilthub.cmu.edu/articles/dataset/Blockchain_Address_Poisoning_Companion_Dataset_/29212703/1 We also provide a Jupyter notebook explaining 1) how to access the dataset, 2) how to produce descriptive statistics such as the number of poisoning transfers, and 3) how to manually verify the payoff transfer on Etherscan (BSCscan).
+This dataset will enable other researchers to validate our results as well as conduct further analysis.
 Acknowledgments
 This research was partially supported by a Sui Foundation Academic Grant, the Carnegie Mellon CyLab’s Secure Blockchain Initiative, and the Nakajima Foundation.
-Appendix A Comparison with Guan and Li’s
-This section compares our detection performance to the most recent work from Guan and Li [10].
-To align with their observation, we only include our measurement results from Nov.
-2022 to Feb.
-2024 on Ethereum.
-Table 6 summarizes the results.
-First, the total amount of payoff transfers is similar, corroborating the validity of both approaches.
+Appendix A Investigating false negatives
+This appendix provides more details about the procedure we use to categorize false negatives in §5.2.
+We first obtain the set of victims $V$ that are seemingly attacked by lookalike addresses $L$ that we had missed in the Guan and Li’s dataset [10] we used for evaluation.
+We then retrieve each victim’s ERC-20 or native ETH transfers on Etherscan.
+We follow the flow chart laid out in Figure 13 and use the criteria defined in the green boxes to determine the reason (in the blue boxes) why we missed the lookalike address.
+“Mislabeling” indicates that the false negatives were actually true negatives – and that the issue lied with the dataset we compared to.
+(There are 45 such cases.) The asterisk (*) indicates that we lower the similarity threshold if we do not find the intended address $R$ with the current threshold $(3,4)$.
+The “unknown” category (orange box) signifies that the reason is (primarily) due to data collection quirks.
+For example, we fail to capture the original transfer between the victim $V$ and the intended address $R$ if the transfer is outside of our measurement period (2 years) or uses transferFrom instead of transfer.
+(We capture transferFrom when used for poisoning transfers.) These concern only 38 false negatives out of 164.
+Figure 13: Flow chart of the identification of false negatives causes.
+Appendix B Additional comparisons with other works
+This appendix provides additional details on the performance comparison of our detection algorithm to Ye et al [47] and Guan and Li [10].
+For a fair comparison, here, we restrict the data we collected to the same time period as those studies.
+We detect more payoff transfers than Ye et al.
+We find 1,059 payoff transfers with 36.6M USD in losses while they found 560 transfers with at most 28M USD in losses, including two non-address poisoning attacks they investigate.
+For poisoning transfers, we identify 4.9M zero-value transfers, compared to 21M for Ye et al.
+It is unclear whether the difference comes from 1) false positives on Ye et al.’s side (i.e., benign zero-value transfers), or 2) false negatives on our side.
+Table 7 summarizes our comparison with Guan and Li [10].
+First, we find a similar total amount of payoff transfers, corroborating the validity of both approaches.
 Second, both approaches estimate similar attack costs in ETH, but not in USD.
-Guan and Li use a fixed ETH to USD conversion rate (around 3200 USD) whereas we query the price at the time of attack.
-The exchange rate, however, fluctuated between 1000 USD and 2000 USD for the majority of the observation period, leading to the attack costs being overestimated by over 10 million USD, which results in a difference of a 300% on the return-on-investment (RoI) for the attacker.
-Third, we detect more than 2.2 million counterfeit token transfers.
-Guan and Li do not report tokens with symbols that do not exactly match USDT and USDC.
-For example, out of 5,838 counterfeit tokens, only 1,922 use the exact same token symbols whereas at least 740 tokens use the different encodings of USDT or USDC (e.g., using a Cyrillic “S” and/or “T” in “USDT”).
-Likewise, attackers can also avoid their detection if they change the attack characteristics, e.g., launching zero-value transfers without attack contracts.
-Our results suggest filtering the transfers based on address similarity first would be more robust to attacker manipulation (less false negatives), than filtering by attack characteristics.
-On the other hand, compared to Guan and Li [10], our technique misses attacks where less than 7 digits match the address $L$.
-In addition to the methodological differences, our work has several distinctions.
-First, we have better coverage in measurement (two years, two chains).
-We find 270M attack attempts (compared to 14M), highlighting significant chain waste and user experience degradation.
+Guan and Li use a fixed ETH-USD conversion rate (around 3,200 USD) whereas we query the price at the time of attack.
+Because the exchange rate fluctuated between 1000 and 2000 USD for the majority of the observation period, Guan and Li overestimate attack costs by over 10 million USD, which results in a difference of a 300% on the return-on-investment (RoI).
+Third, we have more than 2.2 million counterfeit token transfers and 3,500 counterfeit tokens.
+Guan and Li appear to only report counterfeit tokens whose symbols are USDT or USDC.
+For example, out of 5,838 counterfeit tokens we find, only 1,922 use the exact same token symbols whereas at least 740 tokens use the different encodings of USDT or USDC (e.g., using a Cyrillic “S” and/or “T” in “USDT”).
+Finally, our technique misses attacks where $L$ matches $R$ in less than 7 digits.
+Guan and Li do consider these attacks, which, however, are shown to have a negligible impact in §5.2 (and which come at the increased risk of false positives).
 Table 7: Comparison of detection performance
-Payoff Cost ETH ROI Tiny Zero-value Counterfeit CT transfer (USD) token Guan and Li [10] 76.79M 7918.2 (25.5M) 2.2 277.3K 7.85M 6.64M 2,333 Ours 76.81M 8202 (14.9M) 5.1 283.9K 7.05M 8.87M 5,838
-Second, although Guan and Li also perform attacker clustering, they appear to have overestimated group sizes, due to erroneous merges from copying bots, as we show in §5.3.
-We carefully remove bots and extrapolate a few large attack entities.
-Last but not least, most of our analyses are novel.
-We analyze victim characteristics, success conditions, group competitions, cross-chain attacks, and modeling of address generation for the first time.
-We provide deeper insights on how attackers perform attacks and what capabilities they have.
-Appendix B Birthday Paradox
+Payoff Cost ETH ROI Tiny Zero-value Counterfeit CT transfer (USD) token Guan and Li [10] 76.79M 7918.2 (25.5M) 2.2 277.3K 7.85M 6.64M 2,333 This paper 76.81M 8202 (14.9M) 5.1 283.9K 7.05M 8.87M 5,838
+Appendix C Impact of the Birthday Paradox
 This section investigates whether our detection system mislabels poisoning transfers when the victim interacts with two similar-looking addresses by coincidence (i.e., due to the “birthday paradox”).
@@
 For instance, when the victim interacts with $r=$19,290 accounts, the probability $p$ becomes 0.5.
-Here, we examine whether our system should exclude a set of victims that present a collision probability greater than a given threshold $a$, i.e., $p\geq a$.
+We examine whether our system should exclude a set of victims that present a collision probability greater than a given threshold $a$, i.e., $p\geq a$.
 We denote this set of victims by $\mathcal{V}_{p\geq a}$.
-Figure 13(a) shows the number of victims ($|\mathcal{V}_{p>a}|$) that present a collision probability greater than $a$, on the $x$ axis.
+Figure 14 shows the number of victims ($|\mathcal{V}_{p>a}|$) that present a collision probability greater than $a$, on the $x$-axis.
 Even at a small threshold $a=0.01$, we only find 664 accounts (0.04% of total victims).
-At first, excluding $\mathcal{V}_{p\geq a}$ does not seem to impact our detection performance.
-However, we find that excluding these accounts results in numerous false negatives because attackers purposefully target very active accounts, as shown in §6.1.
-Figure 13(b) illustrates the number of attacks we would report if we excluded attacks on $\mathcal{V}_{p>a}$.
+Excluding $\mathcal{V}_{p\geq a}$ does not seem to impact our detection performance.
+However, excluding these accounts results in numerous false negatives because attackers purposefully target very active accounts, as shown in §6.1.
+Figure 14 illustrates the number of attacks we would report if we excluded attacks on $\mathcal{V}_{p>a}$.
 For example, we would discover 16.2M million poisoning transfers at $a=0.01$, which is 1 million fewer than at $a=0.99$.
 Indeed, we find that most of those detected attacks are indeed poisoning transfers.
-99.7% of attacks targeting $\mathcal{V}_{p>0.99}$ are launched by the attack accounts $\mathcal{A}$ that have also targeted victims $\mathcal{V}_{p<0.01}$ (i.e., the population that is hardly susceptible to a collision).
-False positives on payoff transfers are even more unlikely as they require all of the following: 1) the victim interacts with two similar addresses, 2) one of similar addresses “unintentionally” performs the poisoning transfer (e.g., testing with a zero transfer), and 3) the intended transfer, the poisoning transfer, and the payoff transfer happen in that correct order.
-As shown in §5.2, we do not find any false positives for top 30 cases for both chains.
-The result shows that our system is robust against spurious collisions and that setting a low threshold reduces the performance of our detection system.
-Hence, we decide to set the threshold of $a=0.999$ to exclude victims that are interacting with a large enough number of addresses to be extremely susceptible to collisions, while capturing nearly all attacks.
-Setting such a high threshold only excludes, for instance, 64 Ethereum accounts.
-(a) Nr. of accounts ($y$-axis) with a probability of collision greater than a given threshold ($x$-axis)
-(b) Nr. of poisoning transfers detected at a given collision probability threshold
-Figure 13: The number of victims that are susceptible to a collision and the number of attacks that are detected when we exclude those victims. The $x$-axis does not start with 0.
-Appendix C Accidental transfers
+99.7% of attacks targeting $\mathcal{V}_{p>0.99}$ are launched by the same attack accounts $\mathcal{A}$ that have also targeted victims $\mathcal{V}_{p<0.01}$ (i.e., the population that is hardly susceptible to a collision).
+False positives on payoff transfers are even more unlikely as they require: 1) the victim interacts with two similar addresses, 2) one of similar addresses unintentionally performs the benign transfer (e.g., a zero transfer for testing), and 3) the intended, the poisoning, and the payoff transfers happen in that correct order.
+We indeed do not find any false positives for the top 30 cases for both chains in §5.2.
+Those results imply that our system is robust against spurious collisions and that setting a low threshold reduces the performance of our detection system.
+Hence, we decide to set $a=0.999$ to exclude victims that are interacting with a large enough number of addresses to be extremely susceptible to collisions, while capturing nearly all attacks.
+Setting such a high threshold only excludes 64 Ethereum accounts.
+Figure 14: Left: Nr.
+of victims that are susceptible to a collision ($y$-axis) given a threshold ($x$-axis).
+Right: Nr.
+of attacks detected ($y$-axis) when we exclude those victims with a threshold ($x$-axis).
+The $x$-axis does not start with 0.
+Appendix D Accidental transfers
 Victims could accidentally send tokens to an incorrect, but not malicious, destination address, due to a typing mistake.
@@
 “3,” “3” vs.
-“e,” 2) swap the order of two characters, or 3) mistype a character for something visually similar [32]: “b” vs “d,” “e” vs “c,” or “6” and “b.” To quantitatively assess typing mistakes, we calculated the Levenshtein distance $\ell$ (i.e., the number of operations necessary to replace a string with one another, including transposition) between two addresses.
+“e,” 2) swap the order of two characters, or 3) mistype a character for something visually similar [33]: “b” vs “d,” “e” vs “c,” or “6” and “b.” To quantitatively assess typing mistakes, we calculate the Levenshtein distance $\ell$ (i.e., the number of operations necessary to replace a string with one another, including transposition) between two addresses.
 As expected, the distance is small; 84.95% of typos have $\ell=1$, and 96.69% have $\ell\leq 2$ in Ethereum.
 We have similar results in BSC (86.48% and 94.34% for $\ell=1$ and $\ell\leq 2$).
-Appendix D Detecting copying bots
-Here we explain why we use the attack ratio to exclude bots and show the robustness of our clustering results through temporal clustering.
+Appendix E Detecting copying bots
+Here we explain why we use the attack ratio to exclude bots and show robustness in our results via temporal clustering.
 As discussed in §4.2, we group two poisoning transfers if they 1) are in the same $TX$, 2) are launched by the same attacker $A$, or 3) specify the same lookalike address $L$.
@@
 We notice that two seemingly distinct large groups (based on their strategies and behaviors) are connected by a few attack addresses $A$ that behave differently from other attack addresses.
-We find that they appear to be the copying bots that duplicate transactions from multiple attack groups.
+We find that they appear to be the bots that simply copy transactions from multiple attack groups.
 For instance, the account 0x7D575a7C732D1c502c07f18BB822D29CF7DBf9E8 copies not only address poisoning transactions [^24] The original transaction: 0x725eaedf8857e243587020de97e-
@@
 cae7a2ca9bef0e9c1c6202c2ea257c
-but also other transactions such as from MEV (Maximum Extractable Value) bots [^19] The original transaction: 0x0720fec6980c196b3545d273da2-
+but also other transactions such as from MEV (Maximum Extractable Value) bots. [^25] The original transaction: 0x0720fec6980c196b3545d273da2-
 b5b41987bf29524c6236cbf462d683df06057
@@
 87b6732e8b8f27097e8d690c090cf6e6865bf
-. This type of copying behavior is relatively common in Ethereum or BSC, as discussed in previous work [27, 26].
+This type of copying behavior is relatively common in Ethereum or BSC, as discussed in previous work.[28, 27]
 Those bots seem to copy the original transactions without properly understanding address poisoning.
 Without changing the lookalike address $L$, the payoff transfer goes to the original attacker regardless of who initiates the attack.
-In other words, these bots are helping other attackers.
-We next explain why we choose to use the attack ratio to detect copying bots over other methods (using the notion of time or removing all duplicated transactions).
+These bots are helping other attackers.
+We next explain why we use the attack ratio to detect copying bots over other methods (using the notion of time or removing all duplicated transactions).
 Using the notion of time.
@@
 This can be achieved by monitoring unconfirmed transactions in the peer-to-peer (P2P) network.
-For instance, the account 0xfffdcF2B3419C243A5eba5f051A64ad629362c9a manages to include copy transactions in the same block as the original transaction most of the time.
-Based on the public mempool data from Flashbots [^20] https://mempool-dumpster.flashbots.net/index.html, we realize that this account’s transactions do not appear in the public P2P network, while the original transactions do.
-This suggests that some bots monitor the original transactions in the P2P layer, and directly submit their transactions to the block builders (to potentially front-run the original transaction).
+For instance, the account 0xfffdcF2B3419C243A5eba5f051A64ad629362c9a manages to include copy transactions in the same block as the original transaction.
+Based on the public mempool data from Flashbots, [^26] https://mempool-dumpster.flashbots.net/index.html we realize that this account’s transactions do not appear in the public P2P network, while the original transactions do.
+This implies that some bots monitor the original transactions in the P2P layer, and directly submit their transactions to the block builders (to front-run the original transaction).
 As a result, we cannot systematically determine whether the first transaction is the original or the copy.
@@
 This method appears to capture all the transactions from the bots.
-However, we realize that the attackers (not the copying bots) also re-send their poisoning transfers several times against the same victim.
-We would remove more than 11 million (62%) “duplicated” transfers with many false positives, which significantly undermines the final clustering results (i.e., the size of the clusters becomes exceedingly small).
+However, we realize that attackers (not the copying bots) also resend their poisoning transfers several times against the same victim.
+We would remove more than 11 million (62%) “duplicated” transfers, which significantly undermines the final clustering results (i.e., the size of the clusters becomes exceedingly small).
 Removing potential bots by attack ratio.
-Only using information from transactions either leads to 1) the removal of too many non-bot transactions or 2) the inclusion of copying bot transactions.
-We decide to retrieve additional information of attack accounts $A$, and exclude $A$ with a low attack ratio—the ratio of the transactions used for address poisoning, as introduced in §4.2.
+Only using attack transactions either leads to 1) the removal of too many non-bot transactions or 2) the inclusion of copying bot transactions.
+We decide to retrieve additional information of attack accounts $A$, and exclude $A$ with a low attack ratio—the ratio used for address poisoning, as introduced in §4.2.
 We aim to set the threshold as low as possible while removing all the copying bots.
@@
 Therefore, we choose a threshold of 0.5.
-Figure 14: Left: Top 10 clusters ($y$-axis) based on the threshold ($x$-axis). Right: the ratio of removed transfers over the total number of transfers based on the threshold
-Finally, to additionally validate the robustness of our clustering, we perform a temporal clustering.
-We apply our clustering algorithm for the data until time $x$ and gradually increase $x$ over time to observe the change in the clustering results.
+Figure 15: Left: Top 10 clusters ($y$-axis) based on the threshold ($x$-axis). Right: the ratio of removed transfers over the total number of transfers based on the threshold.
+To additionally validate our clustering robustness, we perform a temporal clustering.
+We apply our algorithm for the data until time $x$ and gradually increase $x$ over time to observe the change in the results.
 Figure 16 shows how the top 10 clusters (based on the number of lookalike addresses) evolve over time where the link indicates the same attack group.
-The left figure shows the result with a threshold of 0.5; the clusters appear to be stable over time.
-The right figure is the result without any threshold (no bot exclusion); some large clusters remain separate for a while, but eventually connected around July and August 2023, forming a single large cluster due to the appearance of copying bots.
+The red line denotes Group 1; the blue lines represent other groups.
+The left figure shows results with an attack ratio threshold of 0.5; the clusters appear to be stable over time.
+The right figure shows results without any threshold (i.e., we do not exclude bots); large clusters remain separate for a while, but eventually get connected around Jul.
+and Aug.
+2023, forming a single large cluster due to the emergence of copying bots.
 (a) With a threshold of 0.5
 (b) Without a threshold
-Figure 15: Temporal clustering results. The x-axis is the time, and the y-axis is the number of clusters.
-Appendix E Intended addresses
-We focus here on the intended addresses $R$—the victim’s recipient which attackers try to impersonate.
-Table 7 lists the top ten intended addresses the attackers impersonate the most (based on the number of lookalike addresses $L$ targeting each $R$).
-We label each address based on the third-party Arkham Intel website.
-[^21] https://intel.arkm.com/ We also aggregate the number of attacks targeting each $R$.
-The table shows that many of $R$ are centralized exchange hot addresses.
+Figure 16: Temporal clustering results. $x$-axis is time, $y$-axis is cluster rank of at time $x$. Red: Group 1; blue: other groups. Links are used to denote the same group.
+Appendix F Intended addresses
+In this section, we focus on the intended addresses $R$—the victim’s recipient, which attackers try to impersonate.
+Table 8 lists the top ten intended addresses that the attackers impersonate the most (based on the number of lookalike addresses $L$ targeting each $R$).
+We label each address based on the third-party Arkham Intel website, [^27] https://intel.arkm.com/ and aggregate the number of attacks targeting each $R$.
+Table 8 shows that many of $R$ are centralized exchange hot addresses.
 Since users do not directly interact with those hot wallets, we do not observe any successful cases for those addresses.
 Yet, those exchanges’ deposit addresses frequently interact with those hot wallets, which may explain why the attackers’ (algorithm) capture those intended addresses (as discussed in §6.2).
-Table 7: Top 10 intended addresses
-Intended address $R$ Label Num of $L$ Num of $TR$ 0xA9D1e0…1d3E43 Coinbase Hot Wallet 143,764 18,229 0x28C6c0…f21d60 Binance Hot Wallet 26,373 16,147 0x974CaA…ECC400 Stake.com Hot Wallet 3,254 2,355 0xF7C8dA…38d921 Devout Deposit 5,242 2,014 0xf89d7b…5EaA40 Bybit Hot Wallet 4,556 1,702 0xBFCd86…23d6aD HitBTC Hot Wallet 3,216 1,569 0x731309…AB4D2a BitGet Hot Wallet 7,547 1,477 0x99870D…9AF1e2 HitBTC Deposit 9,700 1,449 0x3CC936…aeCF18 MEXC Hot Wallet 6,603 1,397 0x75e89d…1dcB88 MEXC Hot Wallet 1,629 1,257
-Appendix F UI Design of wallets and chain scanners
+Table 8: Top 10 most targeted intended addresses
+Intended address $R$ Label Num of $L$ Num of $TR$ 0xA9D1e0...1d3E43 Coinbase Hot Wallet 143,764 18,229 0x28C6c0...f21d60 Binance Hot Wallet 26,373 16,147 0x974CaA...ECC400 Stake.com Hot Wallet 3,254 2,355 0xF7C8dA...38d921 Devout Deposit 5,242 2,014 0xf89d7b...5EaA40 Bybit Hot Wallet 4,556 1,702 0xBFCd86...23d6aD HitBTC Hot Wallet 3,216 1,569 0x731309...AB4D2a BitGet Hot Wallet 7,547 1,477 0x99870D...9AF1e2 HitBTC Deposit 9,700 1,449 0x3CC936...aeCF18 MEXC Hot Wallet 6,603 1,397 0x75e89d...1dcB88 MEXC Hot Wallet 1,629 1,257
+Appendix G UI Design of wallets and chain scanners
 This section examines how address poisoning attacks appear in victims’ wallets or chain scanners.
@@
 Specifically, we conduct our experiment on Dec.
-20th, 2024, and used 159,364 gas for three poisoning transfers which consumes about 1% of one block (0.00014% of gas usage for a day).
-We generate a lookalike address $L$ that matches the first six and the last 5 characters to $R$.
-We follow to $AC$ and $CT$ implementations from the attackers who have uploaded their contracts’ source codes on Etherescan and perform three poisoning transfers on our victim account $V$.
-Our performed attack can be found in $V$ 0xB6e84DF1cE401117C450221ccc6EF502cb0e2284.
+20th, 2024, and use 159,364 gas for three poisoning transfers, which consumes about 1% of one block (0.00014% of gas usage for a day).
+We generate a lookalike address $L$ that matches the first six and the last four characters to $R$.
+We follow to $AC$ and $CT$ implementations from the attackers who have uploaded their contracts’ source codes on Etherscan and perform three poisoning transfers on our victim account $V$.
+Our attack can be found in $V=$0xB6e84DF1cE401117C450221ccc6EF502cb0e2284.
 First, we check the victim’s UI on blockchain wallets (as of Jan.
 2nd, 2025).
-Based on the wallet market share [25], we study MetaMask, Trust Wallet, and Phantom.
-Regarding the address abbreviations (first, last), Metamask is (4,4), Trust Wallet is (2, 4), and Phantom is (4, 4).
-Trust Wallet and Phantom only display tiny transfers but not zero-value or counterfeit token transfers.
-Metamask does not appear to show any ERC-20 token transfer history in the first place.
-When initiating the transfers, users no longer have options to select the recipient from “recently used addresses” but manually specify it on their own.
+Based on their wallet market share [26], we study MetaMask, Trust Wallet, and Phantom.
+All three services abbreviate addresses by only showing the first $a$ and last $b$ characters.
+Metamask uses $(a,b)=(4,4)$, Trust Wallet uses (2, 4), and Phantom uses (4, 4).
+Trust Wallet and Phantom only display tiny transfers but hide zero-value or counterfeit token transfers.
+MetaMask does not seem to show any ERC-20 token transfer history in the first place.
+When initiating transfers, users no longer have options to select the recipient from “recently used addresses” but manually specify it on their own.
 The fact suggests that victims may rely more on blockchain scanning services to find the recipients.
-Second, we check the major blockchain scanning services: Etherscan, Ethplorer, OkLink, and Arkm (as of Jan.
-2nd, 2024).
-Table 8 compares the UI of those blockchain scanners (tiny: T, zero-value: Z, and counterfeit token: C).
-The level of address abbreviations significantly differs by each service.
-Etherscan equally shortens the beginning and the end while others significantly abbreviate one side (or no abbreviation at all).
-Since the attackers mostly focus on the first and last characters now, displaying one side might be effective.
+Second, we check four major blockchain scanning services: Etherscan, Ethplorer, OkLink, and Arkham (as of Jan.
+2nd, 2025).
+The level of address abbreviation significantly differs between services.
+Etherscan elides the middle part of the address while keeping the same number of characters at the beginning and at the end, while others significantly abbreviate one side (or do not abbreviate at all).
+Since attackers mostly focus on collision on the first and last characters, displaying one side might be effective.
 Next, we check the poisoning transfers.
-By default, most services (but Ethplorer) successfully hide zero-value transfers while they fail to detect tiny transfers.
-By changing the setting, Etherscan and OkLink additionally show zero-value, and Arkm displays counterfeit token transfers.
-It is generally easier for them to hide tiny than zero-value (i.e., just detecting the value of 0) while they must implement some heuristics to detect tiny or counterfeit token transfers [^22] Etherscan successfully detects all poisoning transfers as of our visit on Jan.
+Most services (but Ethplorer) successfully hide zero-value transfers (by default) but fail to detect tiny transfers.
+By changing settings, Etherscan and OkLink can show zero-value transfers, and Arkham displays counterfeit token transfers.
+It is generally easier for a service to flag zero-value transfers than tiny/counterfeit token transfers (by just observing their value); scanning services must implement more advanced heuristics to detect tiny or counterfeit token transfers [^28] Etherscan successfully detects all poisoning transfers as of our visit on Jan.
 13th, 2025..
-Note that hiding zero-value transfers hinders the usability if users use zero-value transfers for (benign) testing purposes.
-Most services show token icons or give a warning to $CT$ (based on their list of $AC$), which would be helpful for users to distinguish them.
-Table 8: UI comparison of blockchain scanners
-Addr. abbr. Poisoning transfers Token icon Default Setting change Etherscan 8, 9 T C Z Yes Ethplorer No abbr. T, Z, C No OKLink 4, 11 T, C Z Yes Arkm 36, 0 T C Yes
-Appendix G Payoff transfers
-Table 9 presents the 30 largest payoff transfers on the Ethereum blockchain and BSC as examples for readers to investigate.
-These transactions are the final step of the attack, in which victims mistakenly send funds to the attackers.
-We verify the authenticity of these attacks discovered by our algorithm on Etherscan (BSCscan) through the following steps:
-  1. 1.
-    In the transaction page, get the sender (i.e., from) of the Transfer event
-  2. 2.
-    Go to the sender (victim)’s page and examine the Token Transfer (ERC-20) tab.
-  3. 3.
-    Follow the prompt and change the site settings to disable “Ignore Tokens with Poor Reputation” and “Zero-Value Token Transfers” options.
-  4. 4.
-    Find the payoff transfer transaction and verify that there are attack attempts in earlier blocks.
-For example, in the first case of Ethereum, a transfer of 10,000,000 counterfeit tokens from the victim to the (attacker’s) lookalike address took place in block 17,778,047.
-  5. 5.
-    Find the authentic token transfer events from the victim to the intended address that share similar characters to the lookalike address.
-For example, in the first case of Ethereum, a transfer of 10,000,000 USDT (Tether) is sent from the victim to the address 0xa7B4BAC8f0f9692e56750aEFB5f-
-    6cB5516E90570
-Table 9: Largest payoff transfers on Ethereum and Binance Smart Chain.
-USD Phished is the monetary loss in this single transaction, Recipient is the attacker’s lookalike address, Intended recipient is the address that the victim interacted with, Matched is the number of characters matched between the lookalike and the intended recipient, $a+b$ indicates that the lookalike address matched the first $a$ characters and the last $b$ characters of the intended address.
-(a) Ethereum blockchain
-USD Phished Transaction Hash Recipient (attacker) Intended recipient Matched
-20,000,000 0x08255ca0e42a872559437141fa46980e66d907f7668922467d67515b1ebb4b7f 0xa7Bf48…E90570 0xa7B4BA…E90570 3 + 6
-3,800,000 0x2dea87ff34b0d4712d59e582e810d1a8396e469ce7830a61b9521759b28221f1 0x61b697…074c5C 0x61b9d3…174c5c 3 + 5
-3,554,610 0x11ac26acce0620c50731a94b528419771211779903d6917b05f2bd90e4a6476c 0x1cbB23…9B758a 0x1CbB23…9B758a 7 + 6
-2,030,000 0x0911bd6713493e9ab75ef82cc909114218996f0e717b81b71d7bf29fa06e1622 0x74C9bd…60E1cA 0x74C32c…50E1cA 3 + 5
-2,000,000 0x48e591d562a5098527c0de850ba44ce21014726434184f693707f62877a16994 0xbb2EDb…619455 0xBB217c…219455 3 + 5
-2,000,000 0x5ae3147442cb9cba98b8b3dc1548a7c27dab29847ca2c126059ca806ad4479c9 0xC7B14b…B33A8f 0xc7b57d…533a8F 3 + 5
-1,999,000 0x7cae29b7d215ef67fa183d4455760fc6e8698bfb2c5f1744dec4749d0c4b88c7 0xC7B14b…B33A8f 0xc7b57d…533a8F 3 + 5
-1,660,000 0x86d39a9fd0e4078223179e6f59d2651963edb07899efc9558c43a0a338fb40a5 0xEA0d13…CBeb14 0xEa0ddb…FcEb14 4 + 4
-1,499,990 0xf9948a832bfc217451f056fb3dc2df258c6ad525d7c3c184e07eeeee3df7ed09 0x9339e7…1F4126 0x9339CF…3A4126 4 + 4
-1,200,000 0x371f26eb4b08b37d69b280fa5ea06208750c4bc9cf3258950d63c842f9f01154 0xCbA796…34c994 0xCbA796…34c994 8 + 6
-1,045,150 0xa7be1b408b17853c3f9de8cef05371616d154b646e0f7c556feb72dc87ca9676 0x73435A…ca79F7 0x734659…Ca79F7 3 + 6
-1,000,000 0xb97df9feb55c3849da4551f43995ec0f0fd78ff0e3d28f337934dbaba154bc11 0x80D707…4dbeA2 0x80D713…f4beA2 4 + 4
-960,000 0x5b03c7356fe51575bd4fd190cb83841ec7398d8531dd34d21d89fdc0f776afd5 0x946C8e…f98Cd6 0x9462B5…f98CD6 3 + 6
-865,884 0x71eb719e26a31dbb9b05d38b8a36d73537af35e3a5e0d2ccf4bbb6accbd8cd36 0x55B714…Ba9d79 0x55be3c…Ca9d79 3 + 5
-849,996 0xb09f4c9fe09e6bbeb3abfdbed21a0403a2d8d2b6313ed1cbcdbd620a1f748fb2 0xEB4034…84683d 0xEb4034…84683D 8 + 7
-800,158 0xaeb6f1b730e65ddb05c84d71af8c370dead57c4d1aecafc7e2369fde50875ae6 0x0Cee6B…e70EbF 0x0CEE33…3f0EbF 4 + 4
-749,900 0xecd48ffd8d6a09f1ca18e7acbaad4ff4f70116d70fa65dab842a7cd1651df8a0 0xd7F373…7B7403 0xd7F38D…7B7403 4 + 6
-719,000 0x1576beff54a53f0ef0d881a6242edf4539d0964b5b2431f9e1f08fe781b42de3 0x9887E5…D40ea3 0x9887F3…740ea3 4 + 5
-710,000 0x66096743f07f2f0d49818f5e4de28b98bbf3ac6bfcf665fa48225f97260be6b1 0x949D0D…ef861B 0x949954…58861B 3 + 4
-699,990 0xb2d6e74cba922beeae6dfd9f8a832372ae1d4d5314c0d1505dda13911e840316 0xC7BA77…26b438 0xc7BD59…F5B438 3 + 4
-618,000 0x273166dfae3d85e8297a0c1353d8d9bc72ab7b021a61797cd30b8d1871f59fea 0x55Eb2A…A83159 0x55e5Cf…083159 3 + 5
-600,000 0xf997e3b1fcba63c52815dbc63d4364691a79ef65bb578957f182f54753ef59ca 0xE07222…6adFeA 0xE07C8D…6AdFea 3 + 6
-561,035 0x65f005042adff7a66fd5e317459987367dcfbae3246a50dbde88d1f364c4d9a5 0x368595…c3aFd3 0x3683a1…C3aFD3 3 + 7
-541,900 0x82be73572215148ba7c03e984b1361188b098b7ceedf1aaac2a2574ec3f52fa0 0x35D6CD…4B0305 0x35D3c7…4b0305 3 + 6
-541,000 0x7b5b69fad190d0e2b233d1b5bc0fb4a92178b72a737b3f2041a438a8bcd8ecad 0x60855B…30D466 0x6085fd…c6D466 4 + 4
-500,962 0xea71be058fcd162620d95d04b2d5acbec953cedbb5ccb5339e6d43198b4b026e 0x539592…0DDf0b 0x539683…a6Df0B 3 + 4
-500,000 0x8e452c91d99185a0ba325b2ea3dbb276d09f4b2a6dc6772d615acd48237a3148 0x324AFa…80C28F 0x32494C…C0c28F 3 + 5
-500,000 0xd4bc53f71576fc1b78dd56145e42a127d4109567e7d04c3033e59d9ce9da502c 0xB08cfc…A74a2F 0xb08D97…574A2f 3 + 5
-500,000 0xe68134277191eae950ca698b6651ac7c1942c6c4578eb53a98379c7deceb3f33 0x140588…b0Cb01 0x140b09…25cB01 3 + 4
-500,000 0x78cfaa6c26a06444561397b7b02a426b1e3250b1eaf7e051e91b9137a42fdf86 0xC6BCCb…Bf0115 0xc6B7A8…BF0115 3 + 6
-(b) Binance Smart Chain
-USD Phished Transaction Hash Recipient (attacker) Intended recipient Matched
-279,489 0xa0880a04b9f4dd114eb6e5b457a4b6f260c1e1444eef350b62cdc5f1cba8b0ca 0x7e4d7a…a7f223 0x7e4D54…dcf223 4 + 4
-150,000 0xba144f4c4542434bc1ebc2d70030a89f46e81c993842bf79b43ad453cd446d3e 0x112C71…C4B08a 0x112784…24B08a 3 + 5
-144,876 0xc11c5930465a2c33a479b7df63eaa15610dd9ca78c70f788d0f69cfe5e96960d 0x25f566…b02c50 0x25f5e8…b02c50 4 + 6
-133,000 0x266647bc5190cd516f791399ac61b10e876ad75dcc0c6fabb77ccc80f7ead038 0xe72a80…49dc98 0xe72dee…f7dc98 3 + 4
-100,000 0x62798e183339f5ed70239a84b4463b7e48f497e1ef0e23ee82b78087195462d4 0xD643f1…eAF6fB 0xd64329…24F6fB 4 + 4
-97,250 0x9a4081519e99e2e6fe6add185da7333d44eb28cd1c70df36a32ff8648940cbf5 0x8AC739…5cEB7B 0x8ac7CE…76eb7b 4 + 4
-91,680 0x12505feedc126def4e118a2b8e970766795d45201a79dbe45f4b2b9a41391375 0x5ebefa…8De436 0x5EBe2a…EDe436 4 + 5
-86,180 0xa070b46843b57ff4e529cbe7aabb67c99cb36da58510106bb464d1f039e7b3db 0x2556be…757cF1 0x255EC6…A57cF1 3 + 5
-72,671 0x4e7dd743e32d51f6f8d0705d20542a15a9a8b40a80e707ad1e1f18147033b81b 0x6c9351…D01114 0x6c93Ce…EC1114 4 + 4
-70,006 0x13aecc1f514cc95db38a31ada1c9827124aa87360dc11cb1b3dfd7fe0d2499e0 0x2f9233…2BC04D 0x2f9276…cbc04d 4 + 5
-60,000 0x521bb785beaac9b08cc35d57096a40d28dc7e3e697fe22572a0431681770cc04 0x1616bB…4d1800 0x161c2C…6d1800 3 + 5
-50,111 0xe5576fff71817e990959b6947aa54866888953e2912d3047620b27d0ac766350 0xf06e0a…b8b610 0xf06e59…29b610 4 + 4
-50,094 0x09c21db09c5e92e68bf01eb46374796a9a3c16a0b42df9ef2217020f2f8e2352 0x5c3415…82E06f 0x5c3464…a2E06F 4 + 5
-50,000 0x6108401848776037926010c09b848f00ec5e04e3c4075b72437575544e4c0cd9 0x809799…4A2d4c 0x8097dc…ba2D4C 4 + 5
-50,000 0xb4a277c05232537bbb1dd2aa1f04c8dced24fbd5be125c65fe0498db3f4b837d 0x1487c9…f08B32 0x148799…b08b32 4 + 5
-50,000 0xedfe46269453c28ff4d70f736ca73664ca032bbc65a67462b3d35ea178e47171 0xC4d43e…71d48D 0xc4D89C…81D48D 3 + 5
-46,000 0xe4335fecccba5eec21dbaa18d71cdb4d1c56056ee38c8a30b2a57aaccf4c4faa 0xcc3af0…97298f 0xcc3afe…fe298f 5 + 4
-40,008 0x230ef03a64905d8e0c9a074fde33c0fb993ad536b3852129c36efe485409ad69 0x458728…dF9961 0x4587C1…629961 4 + 4
-39,995 0x5ac785bf1bc5273ca3183471bafce3e1a301afc969f3e5eb26e83f249cca70b4 0x73A31C…E5A703 0x73af35…35a703 3 + 5
-37,000 0x41c732fa968e3816c69ffaebc849233b84a863695344f24328891654b1230edf 0x52C990…0c1961 0x52c908…fC1961 4 + 5
-35,000 0xd96aa7bfbd32720c88ecfddaec21af640a65cb858ddb735d23b5d6b1fae2b1bc 0xe6287d…08d0EA 0xe62d77…f8D0ea 3 + 5
-32,000 0x7cee39ef7f56a08b88346938540be29d58249a33b1b53a1166203d7c467e21c4 0xF632D3…294ee0 0xf63078…194EE0 3 + 5
-30,060 0x53eb5af3962cf511e51b6369071494118460cc11e24234a3551478f3415e0a47 0x4245B2…782880 0x424276…782880 3 + 6
-30,000 0x6bc9148d278302517d49f7602f5c8fd0ef516ea8563996910db846ab029e6994 0x8b8ea3…7A8F9F 0x8B808A…5a8f9f 3 + 5
-30,000 0xa9c015c3b41dd449266cb67fb8c514aadc501fac88a1c49e06d6b51b0c0a9953 0x532D78…C1540f 0x532Dc7…71540f 4 + 5
-30,000 0x6462aaf80cc01700133399d6086cc0381f72ae48b52cdd85529fd7ec45d19fc9 0x2FD1D3…5cc8B8 0x2fd1Da…cCC8B8 5 + 5
-30,000 0x6ef40292302c45c1a0a1756b9b431e4b08f02118ca78c12892377c52e5a84796 0xBB3e51…c3aEAF 0xBb3fd3…93aEAF 3 + 5
-29,377 0x74123f8aef47435e94a4c2291df65bdcf1d6735de97e05fb84974c48a207983d 0xE15D59…5bA949 0xE15f47…6ba949 3 + 5
-28,650 0x1e56b900524521ef21e82392ff8992675a58f0af74a8ddf476c96876533cd5c2 0xaF64F8…Cf01d0 0xaF6859…3f01d0 3 + 5
-25,666 0x60b3ca5895303d08fe83288419e805de7cd078b1778ba7ee30ceafdbac170601 0x4A2274…7b57A0 0x4a22aE…bB57a0 4 + 5
+Hiding all zero-value transfers also hinders usability if users use zero-value transfers for (benign) testing purposes.
+Most services show token icons or give a warning to $CT$ (based on their list of $AC$), which should be helpful to users.